### v1

# Point-to-set Principle and Constructive Dimension Faithfulness

Abstract
We introduce a constructive analogue of $\Phi$-dimension, a notion of Hausdorff dimension developed using a restricted class of coverings of a set. A class of coverings $\Phi$ is said to be “faithful” to Hausdorff dimension if the $\Phi$-dimension and Hausdorff dimension coincide for every set.
We prove a Point-to-Set Principle for $\Phi$-dimension, through which we get Point-to-Set Principles for Hausdorff Dimension, continued-fraction dimension and dimension of Cantor Coverings as special cases. Using the Point-to-Set Principle for Cantor coverings and a new technique for the construction of sequences satisfying a certain Kolmogorov complexity condition, we show that the notions of “faithfulness” of Cantor coverings at the Hausdorff and constructive levels are equivalent.
We adapt the result by Albeverio, Ivanenko, Lebid, and Torbin [1] to derive the necessary and sufficient conditions for the constructive dimension faithfulness of the coverings generated by the Cantor series expansion. This condition yields two general classes of representations of reals, one whose constructive dimensions that are equivalent to the constructive Hausdorff dimensions, and another, whose effective dimensions are different from the effective Hausdorff dimensions, completely classifying Cantor series expansions of reals.
1 Introduction
1.1 Faithfulness in dimension
In the study of randomness and information, an important concept is the preservation of randomness across multiple representations of the same object. Martin-Löf randomness, and computable randomness, for example, are preserved among different base-$b$ representations of the same real (see Downey and Hirschfeldt [5], Nies [27], Staiger [31]) and when we convert from the base-$b$ expansion to the continued fraction expansion ([23], [26], [25]).
A quantification of this notion is whether the rate of information, or constructive dimension, is preserved across multiple representations. This rate is studied using notions such as entropy rates and constructive Hausdorff, and packing dimension. Hitchcock and Mayordomo [8] show that all base-$b$ representations constructive dimension is preserved across base-$b$ representations. However, in a recent work, Nandakumar, Akhil, and Vishnoi [24] show that the rate of information is not preserved across all representations. In particular, they show that constructive dimension is not preserved when we convert from base-$b$ representation to continued fraction representation of the same real.
This raises the following question: Under which settings is the effective rate of information - i.e. constructive dimension - preserved when we change representations of the same real?. Since constructive dimension is a constructive analogue of Hausdorff dimension, this question is a constructive analogue of the concept of “faithfulness” of Hausdorff dimension.
A family of covers $\Phi$ is “faithful” if the dimension of every set $S$ defined using covers from $\Phi$, called the Hausdorff $\Phi$-dimension, coincides with the Hausdorff dimension of $S$. This notion is introduced in a work of Besicovitch [3], which shows that the class of dyadic intervals is faithful for Hausdorff dimension. Faithfulness is well-studied as determining the Hausdorff dimension of a set is often a difficult problem, and faithful coverings help simplify the calculation. Rogers and Taylor [29] further develop the idea to show that all covering families generated by comparable net measures are faithful for Hausdorff dimension. This implies that the class of covers generated by the base $b$ expansion of reals for any $b\in\mathbb{N}\setminus\{1\}$ is faithful for Hausdorff dimension. However, not all coverings are faithful for Hausdorff dimension. A natural example is the continued fraction representation, which is not faithful for Hausdorff dimension [28]. Faithfulness of Hausdorff dimension has then been studied in various settings [1], [2], [9], [28].
1.2 Constructive Dimension Faithfulness
In this work, we introduce a constructive analogue of Hausdorff $\Phi$-dimension which we call constructive $\Phi$-dimension. Recall that base-$b$ representations are faithful for classical Hausdorff dimension. Mayordomo and Hitchcock [8] show that base-$b$ representations are faithful for constructive dimension. On the other hand, Nandakumar, Akhil, and Vishnoi’s work shows that the continued fraction expansion, which is not faithful for Hausdorff dimension is also not faithful for constructive dimension [24].
This raises the natural question whether faithfulness with respect to Hausdorff dimension and faithfulness with respect to constructive dimension are equivalent notions. A positive answer to this question implies that Hausdorff dimension faithfulness, a geometric notion, can be studied using the tools from information theory. Conversely, the faithfulness results of Hausdorff dimension can help us understand the settings under which constructive dimension is invariant for every individual real.
In this work, we show that for the most inclusive generalization of base-$b$ expansions under which faithfulness has been studied classically, namely, for classes of coverings generated by the Cantor series expansions, the notions of Hausdorff faithfulness and constructive faithfulness are indeed equivalent. The Cantor series expansion, introduced by Georg Cantor [4], is a natural generalization of the base-$b$ representation. In contrast to the base-$b$ representation where the basis of representations are exponentials with respect to a fixed $b$, that is $\{b^{n}\}_{n\in\mathbb{N}}$, the Cantor series representation $\{n_{k}\}_{k\in\mathbb{N}}$ uses factorials $\{n_{1}\dots n_{k}\}_{k\in\mathbb{N}}$ as the basis. This class is of additional interest as there are Cantor expansions that are faithful as well as non faithful for Hausdorff dimension, depending on the Cantor series representation $\{n_{k}\}_{k\in\mathbb{N}}$ in consideration [1].
To establish our result, the tool we use a $\Phi$-dimensional analogue of the Point-to-Set Principle.
1.3 Point-to-Set Principle and Faithfulness
The Point-to-Set principle introduced by J.Lutz and N.Lutz[13] relates the Hausdorff dimension of a set with the constructive dimensions of points in the set, relative to an oracle. This theorem has been instrumental in answering open questions in classical fractal geometry using the theory of computing (See [14] [19], [18], [17], [13]). Mayordomo, Lutz, and Lutz [15] extend this work to arbitrary separable metric spaces. We note that this generalization does not cover the $\Phi$-dimension framework, as a restricted class of covers is considered in the case of $\Phi$ dimension.
In this work, we first prove the Point-to-Set principle for $\Phi$-dimension (Theorem 3) and show that this generalizes the original Point-to-Set principle. We then develop a combinatorial construction of sequences having Kolmogorov complexities that grow at the same rate as a given sequence relative to any given oracle (Theorem 6). This new combinatorial construction may be of independent interest in the study of randomness. Using these new tools, we show that under the setting of covers generated by Cantor series expansions, the notions of constructive faithfulness and Hausdorff dimension faithfulness are equivalent (Theorem 9). We then adapt the result by Albeverio, Ivanenko, Lebid, and Torbin [1] to derive a loglimit condition for the constructive dimension faithfulness of the coverings generated by the Cantor series expansions (Theorem 11).
Our main results include the following.
  1. 1.
    We introduce the notion of constructive $\Phi$-dimension that subsumes base-b, continued fraction, and Cantor covering dimension. We prove a point-to-set principle for these families of covers. This generalizes the original Point-to-Set Principle and yields new point-to-set principles for the dimensions of continued fractions and Cantor series representations.
  2. 2.
    Using the point-to-set principle, we characterize constructive faithfulness for Cantor series expansions using a log limit condition of the terms appearing in the series. This generalizes the invariance result of constructive dimension under base $b$ representations to all Cantor series expansions that obey this log limit condition. Moreover, it implies that for any Cantor series expansion that does not obey the log limit condition, there are sequences whose Cantor series dimension is different from its constructive dimension.
The recent works of J. Lutz, N. Lutz, Stull, Mayordomo and others study the “point-to-set principle” of how constructive Hausdorff dimension of points may be used to compute the classical Hausdorff dimension of arbitrary sets. In addition to the generalization of this point-to-set principle to $\Phi$-systems, our final result may be viewed as a new point-to-set phenomenon for the notion of “faithfulness”: here, equality of the constructive Cantor series dimensions and constructive dimensions of every point yield equality for the classical Cantor series and Hausdorff dimensions of every set, and conversely.
1.4 Proof techniques
In Section 3, we introduce $\Phi$-dimension and constructive $\Phi$-dimension which generalizes Hausdorff dimension of well-studied systems in the literature, including base-$b$ and continued fractions. To establish that the classical $\Phi$-dimension is well-defined, we use the standard tools from classical fractal geometry (see for example, Falconer [6]). In Section 4, we prove that the point-to-set principle holds for $\Phi$ systems. The proof adapts the technique in J. Lutz and N. Lutz [13] of constructing the minimizing oracle directly.
To establish the characterization of constructive faithfulness of Cantor series expansions, we utilize a new combinatorial technique. In Section 5, we give a new construction technique for sequences whose relativized Kolmogorov complexities vary together in a controlled manner. In Section 6, we carefully utilize this construction to characterize constructive faithfulness of Cantor series expansions. As an additional application of this technique, we provide a new and more direct proof of a result due to Lutz and Weihrauch [16] that the classical Hausdorff dimension of the set of points with effective Hausdorff dimension $s$, is equal to $s$, for $s\in[0,\infty)$.
2 Preliminaries
We use $\Sigma$ to denote the binary alphabet $\{0,1\}$, $\Sigma^{*}$ represents the set of finite binary strings, and $\Sigma^{\infty}$ represents the set of infinite binary sequences. We use $\lvert x\rvert$ to denote the length of a finite string $x\in\Sigma^{*}$. For an infinite string $X=X_{0}X_{1}X_{2}\dots$, we use $K(X\upharpoonleft n)$ to denote the finite string consisting of the first $n$ symbols of $X$. When $n\geq m$ we also use the notation $X[m,n]$ to denote the substring $X_{m}X_{m+1}\dots X_{n}$ of $X\in\Sigma^{\infty}$. The conditional Kolmogorov complexity of $x\in\Sigma^{*}$ with respect to $y\in\Sigma^{*}$ is defined as,
$K(x\mid y)=\min\limits_{\pi\in\Sigma^{*}}\left\{\lvert\pi\rvert\mid U(\pi,y)=x\right\}$
where $U$ is any fixed prefix-free universal Turing machine (for more details see [5], [10], [27], [30]). $K(x)$ denotes the Kolmogorov complexity of $x\in\Sigma^{*}$ which is equal to $K(x\mid\lambda)$ where $\lambda$ is the empty string. The relativized versions of the above notions $K^{A}(x\mid y)$ and $K^{A}(x)$ are defined by giving the universal Turing machine $U$ an oracle access to the sequence $A\in\Sigma^{\infty}$ ($A$ can be equivalently viewed as a subset of $\mathbb{N}$).
For any $b\in\Sigma^{*}$, the cylinder set of $b$, denoted $C_{b}$, is defined by $C_{b}=\{X\in\Sigma^{\infty}\mid b\sqsubset X\}$, where $u\sqsubset v$ denotes $u$ is a proper prefix of $v$. For a set $U\subseteq\mathbb{R}$, we denote $|U|$ to denote the diameter of $U$, that is $|U|=\sup_{x,y\in U}|x-y|$, where the metric is the Euclidean metric.
For some fixed $n\in\mathbb{N}$, let $\mathbb{X}\subseteq\mathbb{R}^{n}$ denote the metric space under consideration. For any $X\in\mathbb{X}$, we use the notation $X\upharpoonleft n$ to denote the string consisting of the first $n$ bits in the binary expansion of $X$.
We use $[N]$ where $N\in\mathbb{N}$ to represent the integers $\{0,1,\dots,N-1\}$. For any finite collection of sets $\{U\}$, we use #$\#(U)$ to denote the number of elements in $U$.
3 Hausdorff and Effective $\Phi$-dimensions
Hausdorff originally developed the idea of Hausdorff dimension using the notion of s-measures of open covers of a set [7] (see also, for example, Falconer [6]).
We define the general notion of Hausdorff $\Phi$-dimension by restricting the class of admissible covers to the family of covers $\Phi$. In this work, we consider the class of coverings $\Phi$ which we call as a family of finitely intersecting covers which satisfy the properties given below.
Definition 1 (Family of finitely intersecting covers $\Phi$).
We consider the space ${\mathbb{X}}\subseteq\mathbb{R}^{n}$ where $n\in\mathbb{N}$ and a countable family of covers $\Phi=\{U_{i}\}_{i\in\mathbb{N}}$, where for each $i\in\mathbb{N}$, $U_{i}\subseteq\mathbb{R}^{n}$ which satisfy the following properties:
  • Density of Rational points: For each $U\in\Phi$, there exists a $q\in\mathbb{Q}^{n}$ such that $q\in U$.
  • Finite intersection: There exists a constant $c\in\mathbb{N}$ such that for any $\{U_{i}\}\subseteq\Phi$ such that for all $i\neq j$, $U_{i}\not\subseteq U_{j}$ and $\cap_{i}U_{i}\neq\emptyset$, we have #($\#(\{U_{i}\})\leq c$.
  • Fineness : Given any $\epsilon>0$, and $x\in{\mathbb{X}}$, there exists a $U\in\Phi$ such that $|U|<\epsilon$ and $x\in U$.
For any family of finitely intersecting covers $\Phi$, we now define the notions of Hausdorff and Constructive Dimensions with the set of covers restricted to $\Phi$. We later observe that this definition covers the following dimensions.
  • Binary dimension over the space generated by binary alphabets $\Sigma^{\infty}$.
  • Continued fraction dimension over the space generated by continued fraction digits $\mathbb{N}^{\infty}$.
  • Dimension of Cantor Coverings over the space generated by Cantor series expansions.
In contrast to Hausdorff dimension where all the admissible coverings of a set are considered, depending on the class $\Phi$, this can lead to a higher value of dimension.
Discussion: Mayordomo [22] gives a definition of Nice covers of a metric space. They then give the definition of constructive dimension on a metric space with a nice cover. We note here that the setting of finitely intersecting covers is incomparable with the setting of nice covers. Our definition does not require the $𝑜𝑣𝑒𝑟c-c$ property of nice covers. Also, the finite intersection property required in our definition is not there in the definition of nice covers.
3.1 Hausdorff $\Phi$-dimension
Definition 2.
Let $\Phi$ be a family of finitely intersecting covers defined over ${\mathbb{X}}$. Given a set $\mathcal{F}\subseteq{\mathbb{X}}$, a collection of sets $\{U_{i}\}_{i\in\mathbb{N}}\subseteq\Phi$ is called a $\delta$-cover of $\mathcal{F}$ using $\Phi$ if for all $i\in\mathbb{N}$, $|U_{i}|\leq\delta$ and $\mathcal{F}\subseteq\bigcup_{i\in\mathbb{N}}{U_{i}}$.
Definition 3.
Given an $\mathcal{F}\subseteq{\mathbb{X}}$, for any $s>0$, we define
$\mathcal{H}^{s}_{\delta}(\mathcal{F},\Phi)=\inf\left\{\sum_{i}|U_{i}|^{s}:\{U_{i}\}_{i\in\mathbb{N}}\text{ is a }\delta\text{-cover of }\mathcal{F}\text{ using }\Phi\right\}.$
As $\delta$ decreases, the set of admissible coverings decreases. Hence $\mathcal{H}^{s}_{\delta}(\mathcal{F},\Phi)$ increases.
Definition 4.
For $s\in(0,\infty)$, the s-dimensional $\Phi$ outer measure of $\mathcal{F}$ is defined as:
$\mathcal{H}^{s}(\mathcal{F},\Phi)=\lim\limits_{\delta\rightarrow 0}\;\mathcal{H}^{s}_{\delta}(\mathcal{F},\Phi).$
Observe that as with the case of classical Hausdorff dimension, for any $t>s$, if $\mathcal{H}^{s}(\mathcal{F},\Phi)<\infty$, then $\mathcal{H}^{t}(\mathcal{F},\Phi)=0$ (see Section 2.2 in [6]).
Finally, we have the following definition of Hausdorff $\Phi$-dimension.
Definition 5 (Albeverio, Torbin[2]).
For any $\mathcal{F}\subset{\mathbb{X}}$, the $\Phi$-dimension of $\mathcal{F}$ is defined as:
${\mathrm{dim}}_{\Phi}(\mathcal{F})=\inf\{s\geq 0:\mathcal{H}^{s}(\mathcal{F},\Phi)=0\}.$
3.2 Effective $\Phi$-dimension
Let $\Phi$ be a family of covers defined over ${\mathbb{X}}$ satisfying Definition 1 that are also computably enumerable. We formulate the notion of constructive $\Phi$-dimension of a point using Kolmogorov Complexity.
Definition 6.
Given an $r\in\mathbb{N}$, and $X\in{\mathbb{X}}$, define
$K_{r}(X,\Phi)=\min_{U\in\Phi}\{K(U)\;|\;X\in U\text{ and }|U|<2^{-r}\}.$
where for $U\in\Phi$, $K(U)$ is defined as
$K(U)=\min\{K(q)\;|\;q\in U\cap\mathbb{Q}^{n}\}.$
For any oracle $A\subseteq\mathbb{N}$, we can similarly define $K_{r}^{A}(X,\Phi)$ by taking $K^{A}(q)$ instead of $K(q)$.
Definition 7.
For any $X\in{\mathbb{X}}$, the constructive $\Phi$-dimension of $X$ is defined as
$\mathrm{cdim}_{\Phi}(X)=\liminf_{r\rightarrow\infty}\frac{K_{r}(X,\Phi)}{r}$
For any $X\in{\mathbb{X}}$ and $A\subseteq\mathbb{N}$, the constructive $\Phi$-dimension of $X$ relative to $A$ is defined as
$\mathrm{cdim}_{\Phi}^{A}(X)=\liminf_{r\rightarrow\infty}\frac{K_{r}^{A}(X,\Phi)}{r}.$
The constructive $\Phi$-dimension of a set is defined as the supremeum of constructive $\Phi$-dimensions of points in the set.
Definition 8.
For any $\mathcal{F}\subseteq{\mathbb{X}}$, the constructive $\Phi$-dimension of $\mathcal{F}$ is defined as
$\mathrm{cdim}_{\Phi}(\mathcal{F})=\sup\limits_{X\in\mathcal{F}}\mathrm{cdim}_{\Phi}(X).$
3.3 Constructive Dimension and Hausdorff Dimension
Mayordomo [21] extends the result by Lutz [11] to give the following Kolmogorov complexity characterization of constructive dimension of sequences in the Cantor space.
Definition 9 (Lutz [11], Mayordomo [21]).
For any $X\in\Sigma^{\infty}$,
$\mathrm{cdim}(X)=\liminf\limits_{n\to\infty}\frac{K(X\upharpoonleft n)}{n}.$
Mayordomo [22] later gave the following Kolmogorov complexity characterization of constructive dimension of points in $\mathbb{R}^{n}$.
Definition 10 (Mayordomo [22]).
For any $X\in\mathbb{R}^{n}$,
$\mathrm{cdim}(X)=\liminf\limits_{r\to\infty}\frac{K_{r}(X)}{r}.$
where $K_{r}(X)=\min\limits_{q\in\mathbb{Q}^{n}}\{K(q):|X-q|<2^{-r}\}$.
We now state some useful properties of constructive dimension. Lutz [12] show that the constructive dimension of a set is always greater than or equal to its Hausdorff dimension.
Lemma 1 (Lutz [12]).
For any $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}(\mathcal{F})\leq\mathrm{cdim}(\mathcal{F})$.
Further, Lutz[12] also show that the constructive dimension of a set is the supremum of the constructive dimensions of points in the set.
Lemma 2 (Lutz [12]).
For any $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\sup\limits_{X\in\mathcal{F}}\mathrm{cdim}(\mathcal{F})$.
The following lemma states that the constructive $\Phi-$dimension of a sequence is always greater than or equal to its constructive dimension.
Lemma 3.
For any class of coverings $\Phi$ , for all $X\in{\mathbb{X}}$, $\mathrm{cdim}_{\Phi}(X)\geq\mathrm{cdim}(X)$.
Proof.
Consider any $a\in\mathbb{Q}$ such that $s>\mathrm{cdim}_{\Phi}(X)$. We have that for infinitely many $r\in\mathbb{N}$, there exists a $U\in\Phi$ with $|U|<2^{-r}$ and a $q\in U$ having $K(q)<sr+o(r)$. Since $|U|<2^{-r}$, we have that $q\in B_{2^{-r}}(X)$. Since this happens for infinitely many $r$, we have that $\mathrm{cdim}(X)\leq s$. ∎
Definition 11 (Dyadic Family of covers).
Consider the space ${\mathbb{X}}=\mathbb{R}^{n}$. The dyadic Family of covers is the set of coverings $B=\bigcup_{r\in\mathbb{N}}\{[\frac{m_{1}}{2^{r}},\frac{m_{1}+1}{2^{r}}]\times\dots\times[\frac{m_{n}}{2^{r}},\frac{m_{n}+1}{2^{r}}]\}_{m_{1}\dots m_{n}\in[2^{r}]}$.
As the Dyadic Family of covers are faithful with respect to Hausdorff dimension [3], we have the following $\Phi$ dimension characterization of Hausdorff dimension. See [6] for the full definition of Hausdorff dimension.
Theorem 1 (Besicovitch [3]).
For all $\mathcal{F}\subseteq\mathbb{R}^{n}$, we have
${\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{B}(\mathcal{F})$
where ${\mathrm{dim}}(\mathcal{F})$ is the Hausdorff dimension of $\mathcal{F}$.
Proof.
For any cover $U_{i}$, there are at most $2^{n}$ covers in $B$, say $\{V_{i}\}_{i=1}^{2^{n}}$ of size at most $2.|U_{i}|$ such that $U_{i}\subseteq\bigcup_{i=1}^{2^{n}}V_{i}$. Therefore we have that for any $\{U_{i}\}$, there exists $\{V_{i}\}$ such that $\sum_{i}|V_{i}|^{s}\leq 2^{s+n}.\sum_{i}|U_{i}|$. ∎
We also show that the set of covers $B$ is faithful for constructive dimension.
Theorem 2.
For all $\mathcal{F}\subseteq\mathbb{R}^{n}$, we have
$\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{B}(\mathcal{F})$
where $\mathrm{cdim}(\mathcal{F})$ is the constructive dimension of $\mathcal{F}$.
Proof.
It suffices to show that for any $X\in{\mathbb{X}}$, $\mathrm{cdim}_{B}(X)\leq\mathrm{cdim}(X)$. Take any $s>\mathrm{cdim}(X)$. For infinitely many $r\in\mathbb{N}$, there exists a point $q\in\mathbb{Q}$ such that $|X-q|<2^{-r}$ and $K(q)<rs+o(r)$. Consider a machine that takes as input a prefix code for $r,q,e$ where $r\in\mathbb{N},q\in Q$ and $e_{i}\in\{-1,0,1\}$ for $i\in\{1,\dots,n\}$ and outputs $p=(\frac{m_{1}+e_{1}}{2^{r}},\dots,\frac{m_{n}+e_{n}}{2^{r}})$ such that $q\in\{[\frac{m_{1}}{2^{r}},\frac{m_{1}+1}{2^{r}}],\dots,[\frac{m_{n}}{2^{r}},\frac{m_{n}+1}{2^{r}}]\}$, where $m_{i}\in[2^{r}]$ for $i\in\{1,\dots,n\}$. We have that $x,p\in U$ for some $U\in B$ and $|x-p|<\sqrt{n}.2^{-r}$ and $K(p)<rs+o(r)$. It follows that $\mathrm{cdim}_{B}(X)\leq s$. ∎
4 Point-to-set principle for $\Phi$-dimension
J. Lutz and N. Lutz introduced the “point-to-set principle” for Hausdorff dimension [13]. They show that the classical Hausdorff dimension of any set $\mathcal{F}\subseteq\mathbb{R}^{n}$ can be characterized in terms of the constructive dimensions of the points in the set, relative to an oracle $A$.
In this work, we introduce the point-to-set principle for $\Phi$ dimension. We characterize Hausdorff $\Phi$ dimension of any set $\mathcal{F}\subseteq{\mathbb{X}}$ in terms of constructive dimensions of elements in the set, relative to an oracle $A$.
Let $\Phi$ be a constructive family of finitely intersecting covers from Definition 1. We first show that Hausdorff dimension of a set is greater than or equal to the supremum of the relative constructive dimensions of elements in the set, over a minimizing oracle $A$. The proof is a direct generalization of the proof in [13].
Lemma 4.
${\mathrm{dim}}_{\Phi}(\mathcal{F})\geq\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(x)$.
Proof.
By definition of $\Phi$-dimension, we have that for any $s\in\mathbb{Q}$ such that $s>{\mathrm{dim}}_{\Phi}(\mathcal{F})$ and $t\in\mathbb{N}$, there exists a sequence of covers $\{{\mathcal{W}}_{i}^{s,t}\}_{i\in\mathbb{N}}$ where each ${\mathcal{W}}_{i}^{s,t}\in\Phi$ such that $|{\mathcal{W}}_{i}^{s,t}|<2^{-t}$ and $\sum_{i\in\mathbb{N}}|{\mathcal{W}}_{i}^{s,t}|^{s}\leq 1$. Also we have that $\mathcal{F}\subseteq\bigcup_{i}{\mathcal{W}}_{i}^{s,t}$. We assume that we are given an oracle access to two functions, $f_{A}:\mathbb{N}\times\mathbb{N}\times\mathbb{Q}\to\mathbb{Q}^{n}$ and $g_{A}:\mathbb{N}\times\mathbb{N}\times\mathbb{N}\times\mathbb{Q}\to\mathbb{Q}$ The first function, where $f_{A}(i,t,s)$ returns a point $u\in{\mathcal{W}}_{i}^{s,t}$. Note that such a rational point exist inside every cover from Definition 1. The second function $g_{A}(i,t,r,s)$ returns a value $d$ such that $|d-|{\mathcal{W}}_{i}^{s,t}||<2^{-r}$. In other words, it returns a $2^{-r}$ rational approximation to $|{\mathcal{W}}_{i}^{s,t}|$. We now show that for any $x\in\mathcal{F}$, and for any $t\in\mathbb{N}$, there exists a $r>t$ such that for some $U\in\Phi$ having $x\in U$ and $|U|<2^{-r}$, we have there exists a $q\in U$ such that $q\in\mathbb{Q}^{n}$ and $K^{A}(q)<rs+o(r)$. We have that $x\in U$ for some $U\in\{{\mathcal{W}}_{i}^{s,t}\}_{i\in\mathbb{N}}$. Let $|U|\in(2^{-r-1},2^{-r}]$ for some $r\in\mathbb{N}$. We also have $r>t$ as $|U|<2^{-t}$. Consider a machine $M$ that takes as an input $\rho\gamma\sigma\tau$ where each of the input comes from a prefix free set of programs. The machine runs the programs to produce $p,t,r,s$. The machine now iterates through $i$ until it finds the $p^{th}$ cylinder $U$ in the set $\{{\mathcal{W}}_{i}^{s,t}\}$ having $g_{A}(i,t,r+3,s)>2^{-r-2}$. Let the original index of $U$ in the set $\{{\mathcal{W}}_{i}^{s,t}\}$ be $k$. Finally the machine outputs $q=f_{A}(k,t,s)$, that is an element from $U$. For any $t\in\mathbb{N}$, since $\sum_{i\in\mathbb{N}}|{\mathcal{W}}_{i}^{s,t}|^{s}\leq 1$, there are fewer than $2^{s(r+3)}$ sets $V$ in $\{{\mathcal{W}}_{i}^{s,t}\}_{i\in\mathbb{N}}$ such that $|V|>2^{-r-3}$. So it follows that there are fewer than $2^{s(r+3)}$ sets $V$ in $\{{\mathcal{W}}_{i}^{s,t}\}_{i\in\mathbb{N}}$ such that $g_{A}(i,t,r+3,s)>2^{-r-2}$. Therefore, it requires fewer than $rs+3s$ bits to describe $p$. Since $t,r,s$ can be described in $o(r)$ bits, we have that $K^{A}(q)<rs+o(r)$. Now, since $q\in U$ such that $x\in U$ and $|U|<2^{-r}$, we have $K_{r}^{A}(x)<rs+o(r)$. From Definition 7, it follows that $\mathrm{cdim}_{\Phi}^{A}(x)\leq s$. ∎
We now show that over any oracle $A$, the Hausdorff dimension of a set is less than or equal to the supremum of the relative constructive dimensions of elements in the set over $A$.
Lemma 5.
${\mathrm{dim}}_{\Phi}(\mathcal{F})\leq\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(X)$
Proof.
Given any $A\subseteq\mathbb{N}$, let $s^{\prime}>s>\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(X)$. Then for every $X\in\mathcal{F}$, we have $\mathrm{cdim}^{A}_{\Phi}(X)\leq s$. Given any $q\in{\mathbb{X}}$ and $k\in\mathbb{N}$, define $B_{k}(q)$ to the collection of maximal sets in $\Phi$ having size less than or equal to $2^{-k}$ containing $q$.
$B_{k}(q)=\{U\in\Phi\;\lvert\;|U|<2^{-k}\text{ and }q\in U\\ \text{ and for all }V\in\Phi\text{ such that }|V|<2^{-k}\text{ and }q\in U,V\subseteq U\}.$
Note that for any $q\in{\mathbb{X}}$, $B_{k}(q)\neq\emptyset$ due to the fineness condition in Definition 1. From the finite intersection property in Definition 1, it follows that #$\#(B_{k}(q))\leq c$. For any $r\in\mathbb{N}$, define the set $Q_{k}=\{q\in\mathbb{Q}^{n}:K^{A}(q)<ks\}$. Now take
${\mathcal{W}}_{r}=\bigcup_{k>r}\bigcup_{q\in Q_{k}}B_{k}(q).$
Now, for every $x\in\mathcal{F}$, $\mathrm{cdim}^{A}_{\Phi}(x)\leq s$, we have that for every $r\in\mathbb{N}$, there exists a $k\geq r$, and there exists a $U\in\Phi$ having $|U|<2^{-k}$ and a $q\in\mathbb{Q}^{n}$ such that $K^{A}(q)<ks$ and $x,q\in U$. We also have that $U\subseteq V$ for some $V\in B_{k}(q)$. Therefore, $\mathcal{F}\subseteq\bigcup\limits_{{\mathcal{W}}\in{\mathcal{W}}_{r}}{\mathcal{W}}$. Also, we have that
$\sum_{W\in{\mathcal{W}}_{r}}|W|^{s^{\prime}}=\sum_{k=r}^{\infty}\sum_{q\in Q_{k}}\sum_{W\in B_{k}(q)}|W|^{s^{\prime}}.$
Since there are fewer than $2^{ks+1}$ $q$’s such that $K^{A}(q)\leq ks$, we have
$\sum_{W\in{\mathcal{W}}_{r}}|W|^{s^{\prime}}$ $\leq\sum_{k=r}^{\infty}c.2^{ks+1}2^{-ks^{\prime}}$
$=2c\sum_{k=r}^{\infty}2^{k(s-s^{\prime})}$
Since $s-s^{\prime}<0$, this sum converges to 0 as $r\to\infty$. So, we have that for any $A\subseteq\mathbb{N}$, ${\mathrm{dim}}(\mathcal{F})\leq s^{\prime}$ for every $s^{\prime}>\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(X)$. ∎
Combining Lemma 4 and Lemma 5, we have the point-to-set principle for $\Phi$ dimension.
Theorem 3.
For all $\mathcal{F}\subseteq{\mathbb{X}}$,
${\mathrm{dim}}_{\Phi}(\mathcal{F})=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(X).$
4.1 Point to Set Principle for constructive dimension
Consider the space $X=\mathbb{R}^{n}$ and the dyadic family of covers from Definition 11. It is routine to verify that this is a family of finitely intersecting covers. Setting $\Phi=B$ and using Theorem 1 and 2, we have the following point-to-set principle relating Hausdorff and Constructive dimensions. Note that this is the same point-to-set Principle in [13].
Theorem 4 (Lutz Lutz [13]).
For all $\mathcal{F}\subseteq{\mathbb{X}}$,
${\mathrm{dim}}(\mathcal{F})=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}(X).$
4.2 Point to Set Principle for Continued Fraction dimension
The notion of constructive continued fraction dimension was introduced by Nandakumar and Vishnoi [26]. The sequence $[a_{1},a_{2},\dots]$ where each $a_{i}\in\mathbb{N}$ is the continued fraction expansion of the number $x=\cfrac{1}{a_{1}+\cfrac{1}{a_{2}+\cdots}}.$ Akhil, Nandakumar and Vishnoi [24] showed that this notion is different from that of constructive dimension. They gave the following Kolmogorov complexity characterization of continued fraction dimension:
Theorem 5.
Given $Y\in\mathbb{N}^{\infty}$, the continued fraction dimension of $Y$ is
$\mathrm{cdim}_{CF}(Y)=\liminf_{n\rightarrow\infty}\frac{K(Y\upharpoonleft n)}{-\log(\mu(Y\upharpoonleft n))}.$
Consider $\Phi$ to be the set of covers generated by the continued fraction cylinders, that is $\Phi=\{\{[[a_{1},a_{2},\dots a_{n}],[a_{1},a_{2},\dots a_{n}+1]]\}_{a_{1}\dots a_{n}\in\mathbb{N}}\}_{n\in\mathbb{N}}$  [^1] Note that when $n$ is odd, the interval is $[a_{1},a_{2},\dots a_{n}+1]<[a_{1},a_{2},\dots a_{n}]$, we omit that here for the sake of simplicity. . It is routine to verify that this is a family of finitely intersecting covers. It follows that for any $Y\in\mathbb{N}^{\infty}$, $\mathrm{cdim}_{CF}(Y)=\mathrm{cdim}_{\Phi}(Y)$. From Theorem 3, we therefore have the following point-to-set principle for Continued fraction dimension.
Corollary 1.
[…]
5 Kolmogorov Complexity Construction
In this section we give a technical construction which is crucial in proving the results in section 6. The lemma says that given an infinite sequence $X$ and an oracle $A$, for any oracle $B$ there exists a sequence $Y$ whose relativized Kolmogorov complexity (of prefixes) with respect to $B$ is similar to the relativized Kolmogorov complexity (of prefixes) of $X$ with respect to $A$.
Theorem 6.
For all $X\in\Sigma^{\infty}$ and $A\in\Sigma^{\infty}$, given a $B\in\Sigma^{\infty}$, there exists a $Y\in\Sigma^{\infty}$ such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$ and $cdim^{B}(Y)=cdim(Y)$.
Proof.
We construct $Y$ in a stage-wise manner. We start with the empty string $\sigma_{0}=\lambda$ and define finite strings $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}\dots$ inductively. Finally we set $Y=\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3}\dots$.
There exist a constant $c_{K}$ such that for sufficiently large n, there exists some $w\in\Sigma^{n}$ with, $K^{B}(w)\geq n-c_{K}$. Let $M$ be any sufficiently large natural number such that $\log M\geq c_{K}$. For any $m\geq 1$, we do the following in stage $m$. If $m=1$ let $s_{1}=0$ and $e_{1}=M^{2}$. Otherwise we set $s_{m}=(M+m-1)^{2}$ and $e_{m}=(M+m)^{2}$.
If $K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])\geq\log(M+m)$, then let $\sigma_{m}=w_{m}z_{m}$ where $w_{m}$ is any $K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])$-length string such that, $K^{B}(w_{m}\mid\sigma_{0}\sigma_{1}\dots\sigma_{m-1})\geq K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])-c_{K}$ and $z_{m}$ is the string consisting of $e_{m}-s_{m}-K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])$ number of zeroes. If $K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])<\log(M+m)$ then we let $\sigma_{m}$ equal to the string consisting of $e_{m}-s_{m}$ number of zeroes. We define $Y=\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3}\dots$. We show that $Y$ satisfies the required properties.
To prove the first property, it is enough to show that
$\left\lvert K^{A}(X\upharpoonleft(M+m-1)^{2})-K^{B}(Y\upharpoonleft(M+m-1)^{2})\right\rvert\leq c(M+m-1)\log(M+m-1)^{2}.$
In the above $c=c_{K}+2+c_{sym}^{A}+c_{sym}^{B}$ where $c_{sym}^{A}$ and $c_{sym}^{B}$ are the constants from the asymptotic error term in the symmetry of information of Kolmogorov complexity with oracles $A$ and $B$ respectively (see [5]). The above statement implies that
$\left\lvert K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)\right\rvert\leq 4c\cdot\sqrt{n}\log n$
which proves the first part. The proof follows by induction on $n$. Consider the case when $m=1$. By the choice of $\sigma_{1}$ and the symmetry of information,
$K^{B}(Y\upharpoonleft M^{2})=K^{B}(w_{1})+K^{B}(z_{1}\mid w_{1})+O(\log M^{2})$
Now,
$\left\lvert K^{A}(X\upharpoonleft M^{2})-K^{B}(Y\upharpoonleft M^{2})\right\rvert\leq c_{K}+\log M^{2}+c_{sym}^{B}\log M^{2}\leq c\log M^{2}$
which proves the base case. Now, assume the statement for $n\geq 1$. Using the symmetry of information,
$K^{B}(Y\upharpoonleft(M+m)^{2})$ $=K^{B}(Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(z_{m}\mid Y\upharpoonleft(M+m-1)^{2},w_{m})+O(\log(M+m)^{2})$
Similarly,
$K^{A}(X\upharpoonleft(M+m)^{2})$ $=K^{A}(X\upharpoonleft(M+m-1)^{2})$
$+K^{A}(X[(M+m-1^{2}):(M+m)^{2}]\mid X\upharpoonleft(M+m-1)^{2})$
$+O(\log(M+m)^{2}).$
Hence, using the induction hypothesis,
$\left\lvert K^{A}(X\upharpoonleft(M+m-1)^{2})-K^{B}(Y\upharpoonleft(M+m-1)^{2})\right\rvert$
$\leq c(M+m-1)\log(M+m-1)^{2}+c_{K}$
$+2\log(M+m)^{2}+c_{sym}^{A}\log(M+m)^{2}+c_{sym}^{B}\log(M+m)^{2}$
$\leq c(M+m-1)\log(M+m)^{2}+(c_{K}+2+c_{sym}^{A}+c_{sym}^{B})\log(M+m)^{2}$
$=c(M+m)\log(M+m)^{2}.$
Now, we prove the second part. It is enough to show that,
$\left\lvert K(Y\upharpoonleft(M+m-1)^{2})-K^{B}(Y\upharpoonleft(M+m-1)^{2})\right\rvert\leq c^{\prime}(M+m-1)\log(M+m-1)^{2}.$
for every $m\geq 1$. Here $c^{\prime}=c_{K}+2+c_{sym}+c^{B}_{sym}$ where $c_{sym}$ is the constant in symmetry of information for non-relativized KC. This implies that
$\left\lvert K(Y\upharpoonleft n)-K^{B}(Y\upharpoonleft n)\right\rvert\leq 4c^{\prime}\cdot\sqrt{n}\log n$
which gives us the required conclusion. Consider the case when $m=1$. Now,
$K(Y\upharpoonleft M^{2})=K(w_{1})+K(z_{1}\mid w_{1})+O(\log M^{2}).$
By the choice of $w_{1}$, it is easy to verify that
$\left\lvert K(w_{1})-K^{B}(w_{1})\right\rvert\leq c_{K}.$
Therefore,
$\left\lvert K(Y\upharpoonleft M^{2})-K^{B}(Y\upharpoonleft M^{2})\right\rvert$ $\leq c_{K}+2\log M^{2}+c^{B}_{sym}\log M^{2}+c_{sym}\log M^{2}$
which proves the base case. Now, assume the statement for $m\geq 1$. We have,
$K^{B}(Y\upharpoonleft(M+m)^{2})$ $=K^{B}(Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(z_{m}\mid Y\upharpoonleft(M+m-1)^{2},w_{m})+O(\log(M+m)^{2}).$
Similarly,
$K(Y\upharpoonleft(M+m)^{2})$ $=K(Y\upharpoonleft(M+m-1)^{2})$
$+K(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})$
$+K(z_{m}\mid Y\upharpoonleft(M+m-1)^{2},w_{m})+O(\log(M+m)^{2})$
By the choice of $w_{m}$,
$\left\lvert K(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})-K^{B}(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})\right\rvert\leq c_{K}$
Therefore using the induction hypothesis,
$\left\lvert K(Y\upharpoonleft(M+m)^{2})-K^{B}(Y\upharpoonleft(M+m)^{2})\right\rvert$
$\leq c^{\prime}(M+m-1)\log(M+m-1)^{2}+c_{K}$
$+2\log(M+m)^{2}+c_{sym}\log(M+m)^{2}+c_{sym}^{B}\log(M+m)^{2}$
$\leq c^{\prime}(M+m-1)\log(M+m)^{2}+(c_{K}+2+c_{sym}+c_{sym}^{B})\log(M+m)^{2}$
$=c^{\prime}(M+m)\log(M+m)^{2}.$
The proof of the second part is thus complete. ∎
6 Equivalence of Faithfulness of Cantor Coverings at Constructive and Hausdorff Levels
In this section, we show that when the class of covers $\Phi$ is generated by computable Cantor series expansions, the faithfulness at the Hausdorff and constructive levels are equivalent notions.
6.1 Faithfulness of Family of Coverings
We will first see the definition of Hausdorff dimension faithfulness. We then introduce the corresponding notion at the effective level, which we call constructive dimension faithfulness.
A family of covers $\Phi$ is said to be faithful with respect to Hausdorff dimension if the $\Phi$ dimension of every set in the space is the same as its Hausdorff dimension.
Definition 12.
A family of covers $\Phi$ over the space ${\mathbb{X}}$ is faithful with respect to Hausdorff dimension if for all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}_{\Phi}(\mathcal{F})={\mathrm{dim}}(\mathcal{F})$.
We extend the definition to the constructive level as well. A constructive family of covers $\Phi$ is defined to be faithful with respect to constructive dimension if the constructive $\Phi$ dimension of every set is the same as its constructive dimension.
Definition 13.
A constructive family of covers $\Phi$ is said to be faithful with respect to Constructive dimension if for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}_{\Phi}(\mathcal{F})=\mathrm{cdim}(\mathcal{F})$.
The following lemma follows from Definition 8 and Lemma 2. It states that constructive dimension faithfulness can be equivalently stated in terms of preservation dimensions of points in the set.
Lemma 6.
A constructive family of covers $\Phi$ is faithful with respect to Constructive dimension if and only if for all $X\in{\mathbb{X}}$, $\mathrm{cdim}_{\Phi}(X)=\mathrm{cdim}(X)$.
Intervals generated by base-$b$ expansions of a number are faithful with respect to Hausdorff dimension [3] [29] and constructive dimension [8] [31]. On the other hand the class of coverings generated by the Continued fraction expansions of a number is not faithful for either constructive dimension [24] or Hausdorff dimension [28]. This raises the question whether faithfulness with respect to Hausdorff dimension and faithfulness with respect to constructive dimension are equivalent notions.
We investigate the above question from the context of classes of coverings generated by Cantor series expansions. Such class of coverings are called as Cantor coverings. This class is of interest as there are Cantor coverings that are faithful as well as non faithful for Hausdorff dimension, depending on the expansion in consideration [1].
Using the Point-to-Set Principle and properties of Kolmogorov complexity, we show that the notions of faithfulness for Cantor coverings at Hausdorff and Constructive levels are equivalent.
6.2 Cantor coverings over unit interval
We consider the faithfulness of family of coverings generated by the computable Cantor series expansion [4]. Note that the Cantor series expansion is a generalization of the base-$b$ representation, which is the special case when $n_{k}=b$ for all $k\in\mathbb{N}$.
Given a sequence $Q=\{n_{k}\}_{k\in\mathbb{N}}$ with $n_{k}\in\mathbb{N}\setminus\{1\}$, the expression
$x=\sum_{k=1}^{\infty}\frac{\alpha_{k}}{n_{1}.n_{2}\dots n_{k}}$
where $\alpha_{k}\in[n_{k}]$ is called the cantor series expansion of the real number $x\in[0,1]$.
An interesting example is the Cantor series expansion of $e$-2 over $Q=\{k+1\}_{k\in\mathbb{N}}$ [20].
$e-2=\frac{1}{2}+\frac{1}{2.3}+\frac{1}{2.3.4}\dots$
Definition 14.
The class of Cantor coverings $\Phi$ over the space ${\mathbb{X}}=[0,1]$ is the set of covers generated by the Cantor series expansion of $\{n_{k}\}_{k\in\mathbb{N}}$ which is the set of intervals $\{\{[\frac{m}{n_{1}.n_{2}\dots n_{k}},\frac{m+1}{n_{1}.n_{2}\dots n_{k}}]\}_{m\in[n_{1}.n_{2}\dots n_{k}]}\}_{k\in\mathbb{N}}$.
Consider $\Phi$ to be the set of covers generated by a computable Cantor series expansion $Q=\{n_{k}\}_{k\in\mathbb{N}}$. It is routine to verify that this is a family of finitely intersecting covers. From Theorem 3, we therefore have the following point-to-set principle for dimension of Cantor coverings.
Theorem 7.
For all $\mathcal{F}\subseteq{\mathbb{X}}$ and for all computable Cantor coverings $\Phi$, ${\mathrm{dim}}_{\Phi}(\mathcal{F})=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(X).$
6.3 Kolmogorov Complexity Characterization of Cantor Series Dimension
We first show a Kolmogorov Complexity characterization of $\Phi$ dimension for all Cantor coverings $\Phi$.
Theorem 8.
For any $x\in{\mathbb{X}}$, and any computable Cantor covering $\Phi$ generated by $\{n_{k}\}_{k\in\mathbb{N}}$,
$\mathrm{cdim}_{\Phi}(x)=\liminf\limits_{k\to\infty}\frac{K(X\upharpoonleft m_{k})}{m_{k}}$
where $X$ is the binary expansion of $x$ and $m_{k}=\left\lfloor\log_{2}(n_{1}.n_{2}\dots n_{k})\right\rfloor.$
Proof.
For any $s>\mathrm{cdim}_{\Phi}(x)$, we have that for infinitely many $r\in\mathbb{N}$, there exists a $q\in\mathbb{Q}$ and a $U\in\Phi$ having $|U|<2^{-r}$ , such that $x,q\in U$ and $K(q)\leq rs+o(r)$. Let $U=[\frac{i}{n_{1}\dots n_{k}},\frac{i+1}{n_{1}\dots n_{k}}]$ for some $k\in\mathbb{N}$ where $i\in[n_{1}\dots n_{k}]$. From the definition of $m_{k}$, we have that $\frac{1}{2.2^{m_{k}}}<|U|\leq\frac{1}{2^{m_{k}}}$. Consider a machine $M$ that takes in the input $\tau\kappa\delta$, each coming from a prefix free set of programs and produces $q,m_{k},e$ where $q\in\mathbb{Q},m_{k}\in\mathbb{N},e\in\{-1,0,1\}$. The machine outputs $\frac{j+e}{2^{m_{k}}}$ such that $q\in[\frac{j}{2^{m_{k}}},\frac{j+1}{2^{m_{k}}})$ where $j\in\{0,1\dots 2^{m_{k}}\}$. Since $|U|\leq\frac{1}{2^{m_{k}}}$, we have that $x\in[\frac{j+e}{2^{m_{k}}},\frac{j+e+1}{2^{m_{k}}})$ for some $e\in\{-1,0,1\}$, and therefore $X\upharpoonleft m_{k}=\frac{j+e}{2^{m_{k}}}$. Since $|U|>\frac{1}{2.2^{m_{k}}}$ and $|U|<2^{-r}$, we have $m_{k}+1>r$. Since $K(q)<rs+o(r)$, we have $K(q)\leq s.m_{k}+o(m_{k})$ and so we have $K(X\upharpoonleft m_{k})\leq s.m_{k}+o(m_{k})$. Since this happens for infinitely many $k$, corresponding to each $r$, we have that $\liminf\limits_{k\to\infty}\frac{K(X\upharpoonleft m_{k})}{m_{k}}\leq s$. Now consider any $s>\liminf\limits_{k\to\infty}\frac{K(X\upharpoonleft m_{k})}{m_{k}}$, we have that for infinitely many $k$, $K(X\upharpoonleft m_{k})\leq s.m_{k}+o(m_{k})$. Consider a machine that takes as the input $\tau\kappa\delta$, each coming from a prefix free set of programs and produces $q\in\mathbb{Q}$, $m_{k}\in\mathbb{N}$ and $e\in\{0,1,2\}$. The program then finds an $i$ such that $q=\frac{i}{2^{m_{k}}}$ and finally outputs $\frac{2i+e}{2.2^{m_{k}}}$. Let $q=X\upharpoonleft m_{k}=\frac{i}{2^{m_{k}}}$. From the definition of $m_{k}$, we have that $\frac{1}{2.2^{m_{k}}}<|U|\leq\frac{1}{2^{m_{k}}}$ for some $U\in\Phi$ such that $x\in U$. As $|U|>\frac{1}{2.2^{m_{k}}}$, it follows that $p=\frac{2i+e}{2.2^{m_{k}}}\in U$ for some $e\in\{0,1,2\}$. Since $K(q)\leq s.m_{k}+o(m_{k})$, we have that $K(p)\leq s.m_{k}+o(m_{k})$. Therefore, $\mathrm{cdim}_{\Phi}(x)\leq s$. ∎
Definition 9 ensures that when the Kolmogorov complexities of any two $X,Y\in\Sigma^{\infty}$ align over all finite prefixes, their constructive dimensions become equal. From Theorem 8, we get that when this happens, the constructive $\Phi$ dimensions also become equal.
Lemma 7.
For any $X,Y\in{\mathbb{X}}$, $A,B\subseteq\mathbb{N}$ and any class of Cantor coverings $\Phi$, if for all n, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$, then $\mathrm{cdim}^{A}(X)=\mathrm{cdim}^{B}(Y)$ and $\mathrm{cdim}^{A}_{\Phi}(X)=\mathrm{cdim}^{B}_{\Phi}(Y)$.
6.4 Faithfulness of Cantor Coverings
We first show that if $\Phi$ is faithful with respect to constructive dimension, then $\Phi$ is also faithful with respect to Hausdorff dimension.
Lemma 8.
For any class of Cantor coverings $\Phi$, if for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$, then for all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{\Phi}(\mathcal{F})$.
Proof.
We first show that if $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$ for all $\mathcal{F}\subseteq{\mathbb{X}}$ , then for all $A\subseteq\mathbb{N}$ and $X\in[0,1]\setminus\mathbb{Q}$, $\mathrm{cdim}^{A}(X)=\mathrm{cdim}^{A}_{\Phi}(X)$.
Let $B=\emptyset$. From Theorem 6, we have that for all $X\in[0,1]$, and $A\subseteq\mathbb{N}$, there exists a $Y\in[0,1]$, such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K(Y\upharpoonleft n)|=o(n)$. Therefore from Lemma 7, we have that $\mathrm{cdim}^{A}(X)=\mathrm{cdim}(Y)$ and $\mathrm{cdim}^{A}_{\Phi}(X)=\mathrm{cdim}_{\Phi}(Y)$. Since $\Phi$ is faithful with respect to constructive dimension, we have that $\mathrm{cdim}(Y)=\mathrm{cdim}_{\Phi}(Y)$. Therefore we have that $\mathrm{cdim}^{A}(X)=\mathrm{cdim}_{\Phi}^{A}(X)$. Let $\mathcal{F}\subset\mathbb{X}$ be arbitrary. From the point-to-set principle for dimension of Cantor Coverings (Theorem 7), we have that
${\mathrm{dim}}_{\Phi}(\mathcal{F})$ $=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(X)=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{X\in\mathcal{F}}\;\mathrm{cdim}^{A}(X)={\mathrm{dim}}(\mathcal{F}).$
The last equality follows from the point-to-set principle for Hausdorff dimension (Theorem 4). ∎
To prove the converse, we require the construction of set $\mathcal{I}_{s}$ that contains all points in ${\mathbb{X}}$ having constructive dimension equal to $s$.
Definition 15.
Given $s\in[0,\infty)$, define $\mathcal{I}_{s}=\{X\in{\mathbb{X}}\;|\;\mathrm{cdim}(X)=s\}.$
Lutz and Weihrauch [16] showed that the Hausdorff dimension of $\mathcal{I}_{s}$ is equal to $s$. We provide a simple alternate proof of this using the point-to-set principle.
Lemma 9.
Lutz and Weihrauch [16]
${\mathrm{dim}}(\mathcal{I}_{s})=s.$
Proof.
Since for all $X\in\mathcal{I}_{s}$, by definition $cdim(X)=s$, from Lemma 2, it follows that $\mathrm{cdim}(\mathcal{I}_{s})=s$. From Lemma 1, we have that ${\mathrm{dim}}(\mathcal{I}_{s})\leq\mathrm{cdim}(\mathcal{I}_{s})$, from which it follows that ${\mathrm{dim}}(\mathcal{I}_{s})\leq s$. Consider any $X\in\mathcal{I}_{s}$. We have that $\mathrm{cdim}(X)=s$. Let $A=\emptyset$. From Theorem 6, we have that for all $B\subseteq\mathbb{N}$, there exists a $Y\in[0,1]$ such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$ having $\mathrm{cdim}^{B}(Y)=\mathrm{cdim}(Y)$. From Lemma 7, we have that $\mathrm{cdim}^{B}(Y)=\mathrm{cdim}^{A}(X)=\mathrm{cdim}(X)=s$. Therefore, we have that for all $B\subseteq\mathbb{N}$, there exits a $Y\in{\mathbb{X}}$ such that $\mathrm{cdim}(Y)=\mathrm{cdim}^{B}(Y)=s$. Hence $Y\in\mathcal{I}_{s}$ and therefore from the point-to-set principle for Hausdorff dimension (Theorem 4), we have that ${\mathrm{dim}}(\mathcal{I}_{s})=\min\limits_{B\subseteq\mathbb{N}}\sup\limits_{Y\in\mathcal{I}_{s}}\mathrm{cdim}^{B}(Y)\geq s$. ∎
We now show that if $\Phi$ is faithful with respect to Hausdorff dimension, then $\Phi$ would also be faithful with respect to constructive dimension.
Lemma 10.
For any class of Cantor coverings $\Phi$, if for all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{\Phi}(\mathcal{F})$ , then for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$.
Proof.
Given an $X\in{\mathbb{X}}$ having $\mathrm{cdim}(X)=s$, consider the set $\mathcal{I}_{s}$ from Definition 15. From Lemma 9, we have ${\mathrm{dim}}(\mathcal{I}_{s})=s$. Since $\Phi$ is faithful with respect to Hausdorff dimension, we have that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})=s$. We now go on to show that for any $X\in{\mathbb{X}}$ having $\mathrm{cdim}=s$ ( or equivalently, for any $x\in\mathcal{I}_{s}$), $\mathrm{cdim}_{\Phi}(X)\leq s$. This along with Lemma 3 shows that $\mathrm{cdim}(X)=\mathrm{cdim}_{\Phi}(X)$. From Lemma 2 and Lemma 6, we have that for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$. We first show that for all $B\subseteq\mathbb{N}$, there exists a $Y\in\mathcal{I}_{s}$ having $\mathrm{cdim}^{B}_{\Phi}(Y)=\mathrm{cdim}_{\Phi}(X)$. Let $A=\emptyset$. From Theorem 6, we have that for all $X\in[0,1]$, and $B\subseteq\mathbb{N}$, there exists a $Y\in[0,1]$, such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$. Setting $A=\emptyset$ and using Lemma 7, we have that $\mathrm{cdim}_{\Phi}(X)=\mathrm{cdim}^{B}_{\Phi}(Y)$. We now show that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})\geq\mathrm{cdim}_{\Phi}(X)$. From the point-to-set principle for dimension of Cantor Coverings (Theorem 7), we have that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})=\min\limits_{B\subseteq\mathbb{N}}\;\sup\limits_{Y\in\mathcal{I}_{s}}\;\mathrm{cdim}^{B}_{\Phi}(Y)$. From the argument given above we have that for all $B\subseteq\mathbb{N}$, there exists a $Y\in\mathcal{I}_{s}$ having $\mathrm{cdim}^{B}_{\Phi}(Y)=\mathrm{cdim}_{\Phi}(X)$. So it follows that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})\geq\mathrm{cdim}_{\Phi}(X)$. Since we have already shown that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})=s$, it follows that $\mathrm{cdim}_{\Phi}(X)\leq s$. ∎
Therefore, we have the following theorem which states that for the classes of Cantor coverings $\Phi$, faithfulness with respect to Hausdorff and Constructive dimensions are equivalent notions.
Theorem 9.
For any class of Cantor coverings $\Phi$,
$\forall\mathcal{F}\subseteq{\mathbb{X}}\;;\;{\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{\Phi}(\mathcal{F})\iff\forall\mathcal{F}\subseteq{\mathbb{X}}\;;\;\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F}).$
6.5 Log limit condition for faithfulness
Albeverio, Ivanenko, Lebid and Torbin [1] showed that the Cantor coverings $\Phi$ generated by the Cantor series expansions of $Q=\{n_{k}\}_{k\in\mathbb{N}}$ can be faithful as well as non faithful with respect to Hausdorff dimension depending on $Q$. Interestingly, they showed that the Hausdorff dimension faithfulness of Cantor coverings can be determined using the terms $n_{k}$ in $Q$.
Theorem 10 (Albeverio, Ivanenko, Lebid and Torbin [1]).
A family of Cantor coverings $\Phi$ generated by $Q=\{n_{k}\}_{k\in\mathbb{N}}$ is faithful with respect to Hausdorff dimension if and only if
$\lim_{k\rightarrow\infty}\frac{\ln n_{k}}{\ln n_{1}.n_{2}\dots n_{k-1}}=0.$
Using the above result and the result that faithfulness at the constructive level is equivalent to faithfulness with respect to Hausdorff dimension (Theorem 9), we have that the condition stated above provides the necessary and sufficient conditions for Cantor series coverings to be faithful for constructive dimension.
Theorem 11.
A family of Cantor coverings $\Phi$ generated by $Q=\{n_{k}\}_{k\in\mathbb{N}}$ is faithful with respect to Constructive dimension if and only if
$\lim_{k\rightarrow\infty}\frac{\ln n_{k}}{\ln n_{1}.n_{2}\dots n_{k-1}}=0.$ (1)
The Cantor series expansion is a generalization of the base-$b$ representation, which is the special case when $n_{k}=b$ for all $k\in\mathbb{N}$. That is $Q_{b}=\{b\}_{n\in\mathbb{N}}$. Since the condition in Theorem 11 is satisfied by $Q_{b}$ for any $b\in\mathbb{N}$, we have the following result by Hitchcock and Mayordomo about the base invariance of constructive dimension.
Corollary 2 (Hitchcock and Mayordomo [8]).
For any $X\in{\mathbb{X}}$ and $k,l\in\mathbb{N}\setminus\{1\}$, $\mathrm{cdim}_{(k)}(X)=\mathrm{cdim}_{(l)}(X).$ where $\mathrm{cdim}_{(k)}(X)$ represents the constructive dimension of $X$ with respect to its base-$k$ representation.
Note that condition (1) classifies the Cantor series expansions on the basis of constructive dimension faithfulness. As an example, even when $n_{k}={2^{k}}$, condition (1) holds, and therefore $Q=\{2^{k}\}_{k\in\mathbb{N}}$ is faithful for constructive dimension. However, when $n_{k}={2^{2^{k}}}$, condition (1) does not hold, and therefore $Q=\{2^{2^{k}}\}_{k\in\mathbb{N}}$ is not faithful for constructive dimension.
7 Conclusion and Open Problems
We developed a constructive analogue of $\Phi$-dimension and proved a Point-to-Set principle for $\Phi$-dimension. Using this, we show that for Cantor series representations, constructive dimension faithfulness and Hausdorff dimension are equivalent notions. We also provide a loglimit condition for faithfulness of Cantor series expansions.
The following are some problems that remain open
  1. 1.
    Are the faithfulness at constructive and Hausdorff levels equivalent for any system of coverings $\Phi$
  2. 2.
    What is the packing dimension analogue of faithfulness, is there any relationship between faithfulness of Hausdorff dimension, packing dimension, constructive dimension and constructive strong dimension ?