### v2

# Point-to-set Principle and Constructive Dimension Faithfulness

Abstract
Hausdorff $\Phi$-dimension is a notion of Hausdorff dimension developed using a restricted class of coverings of a set. We introduce a constructive analogue of $\Phi$-dimension using the notion of constructive $\Phi$-$s$-supergales. We prove a Point-to-Set Principle for $\Phi$-dimension, through which we get Point-to-Set Principles for Hausdorff dimension, continued-fraction dimension and dimension of Cantor coverings as special cases. We also provide a Kolmogorov complexity characterization of constructive $\Phi$-dimension.
A class of covering sets $\Phi$ is said to be “faithful” to Hausdorff dimension if the $\Phi$-dimension and Hausdorff dimension coincide for every set. Similarly, $\Phi$ is said to be “faithful” to constructive dimension if the constructive $\Phi$-dimension and constructive dimension coincide for every set. Using the Point-to-Set Principle for Cantor coverings and a new technique for the construction of sequences satisfying a certain Kolmogorov complexity condition, we show that the notions of “faithfulness” of Cantor coverings at the Hausdorff and constructive levels are equivalent.
We adapt the result by Albeverio, Ivanenko, Lebid, and Torbin [1] to derive the necessary and sufficient conditions for the constructive dimension faithfulness of the coverings generated by the Cantor series expansion, based on the terms of the expansion.
1 Introduction
1.1 Faithfulness in dimension
In the study of randomness and information, an important concept is the preservation of randomness across multiple representations of the same object. Martin-Löf randomness, and computable randomness, for example, are preserved among different base-$b$ representations of the same real (see Downey and Hirschfeldt [5], Nies [28], Staiger [32]) and when we convert from the base-$b$ expansion to the continued fraction expansion ([24], [27], [26]).
A quantification of this notion is whether the rate of information is preserved across multiple representations. This rate is studied using a constructive analogue of Hausdorff dimension called Constructive dimension [12][22]. Hitchcock and Mayordomo [9] show that constructive dimension is preserved across base-$b$ representations. However, in a recent work, Akhil, Nandakumar and Vishnoi [25] show that the rate of information is not preserved across all representations. In particular, they show that constructive dimension is not preserved when we convert from base-$b$ representation to continued fraction representation of the same real.
This raises the following question: Under which settings is the effective rate of information - i.e. constructive dimension - preserved when we change representations of the same real?. Since constructive dimension is a constructive analogue of Hausdorff dimension, this question is a constructive analogue of the concept of “faithfulness” of Hausdorff dimension.
A family of covering sets $\Phi$ is “faithful” to Hausdorff dimension if the dimension of every set $\mathcal{F}$ defined using covers constructed using $\Phi$, called the Hausdorff $\Phi$-dimension, coincides with the Hausdorff dimension of $\mathcal{F}$. Faithfulness is well-studied as determining the Hausdorff dimension of a set is often a difficult problem, and faithful coverings help simplify the calculation. This notion is introduced in a work of Besicovitch [3], which shows that the class of dyadic intervals is faithful for Hausdorff dimension. Rogers and Taylor [30] further develop the idea to show that all covering families generated by comparable net measures are faithful for Hausdorff dimension. This implies that the class of covers generated by the base $b$ expansion of reals for any $b\in\mathbb{N}\setminus\{1\}$ is faithful for Hausdorff dimension. However, not all coverings are faithful for Hausdorff dimension. A natural example is the continued fraction representation, which is not faithful for Hausdorff dimension [29]. Faithfulness of Hausdorff dimension has then been studied in various settings [1], [2], [10], [29].
1.2 Constructive Dimension Faithfulness
In this work, we introduce a constructive analogue of Hausdorff $\Phi$-dimension which we call constructive $\Phi$-dimension. A family of covering sets $\Phi$ is “faithful” to constructive dimension if the constructive $\Phi$-dimension of every set $\mathcal{F}$ coincides with the constructive dimension of $\mathcal{F}$. Mayordomo and Hitchcock [9] show that all base-$b$ representations of reals, which are faithful for Hausdorff dimension, are also faithful for constructive dimension. On the other hand, Nandakumar, Akhil, and Vishnoi’s work shows that the continued fraction expansion, which is not faithful for Hausdorff dimension is also not faithful for constructive dimension [25]. This raises the natural question: Are faithfulness with respect to Hausdorff dimension and faithfulness with respect to constructive dimension equivalent notions ? A positive answer to this question implies that Hausdorff dimension faithfulness, a geometric notion, can be studied using the tools from information theory. Conversely, the faithfulness results of Hausdorff dimension can help us understand the settings under which constructive dimension is invariant for every individual real.
In this work, we show that for the most inclusive generalization of base-$b$ expansions under which faithfulness has been studied classically, namely, for classes of coverings generated by the Cantor series expansions, the notions of Hausdorff faithfulness and constructive faithfulness are indeed equivalent. The Cantor series expansion, introduced by Georg Cantor [4], uses a sequence of natural numbers $Q=\{n_{k}\}_{k\in\mathbb{N}}$ as the terms of representation. Whereas base-$b$ representation use exponentials with respect to a fixed $b$, $\{b^{n}\}_{n\in\mathbb{N}}$, the Cantor series representation $Q=\{n_{k}\}_{k\in\mathbb{N}}$ uses factorials $\{n_{1}\dots n_{k}\}_{k\in\mathbb{N}}$ as the basis for representation. This class is of additional interest as there are Cantor expansions that are faithful as well as non faithful for Hausdorff dimension, depending on the Cantor series representation $\{n_{k}\}_{k\in\mathbb{N}}$ in consideration [1]. To establish our result, we use a $\Phi$-dimensional analogue of the Point-to-Set Principle.
1.3 Point-to-Set Principle and Faithfulness
The Point-to-Set principle introduced by J.Lutz and N.Lutz[14] relates the Hausdorff dimension of a set of $n$-dimensional reals with the constructive dimensions of points in the set, relative to a minimizing oracle. This theorem has been instrumental in answering open questions in classical fractal geometry using the theory of computing (See [15], [21], [20], [19], [14]). Mayordomo, Lutz, and Lutz [16] extend this work to arbitrary separable metric spaces.
In this work, we first prove the Point-to-Set principle for $\Phi$-dimension (Theorem 3) and show that this generalizes the original Point-to-Set principle. We then develop a combinatorial construction of sequences having Kolmogorov complexities that grow at the same rate as a given sequence relative to any given oracle (Theorem 5). This new combinatorial construction may be of independent interest in the study of randomness. Using these new tools, we show that under the setting of covers generated by Cantor series expansions, the notions of constructive faithfulness and Hausdorff dimension faithfulness are equivalent (Theorem 7). We then adapt the result by Albeverio, Ivanenko, Lebid, and Torbin [1] to derive a loglimit condition for the constructive dimension faithfulness of the coverings generated by the Cantor series expansions (Theorem 9).
Our main results include the following.
  1. 1.
    We introduce the notion of constructive $\Phi$-dimension using that subsumes base-b, continued fraction, and Cantor covering dimension. We also give an equivalent Kolmogorov Complexity characterization of constructive $\Phi$-dimension. We prove a Point-to-Set principle for $\Phi$-dimension. This generalizes the original Point-to-Set Principle and yields new Point-to-Set principles for the dimensions of continued fractions and Cantor series representations.
  2. 2.
    Using the point-to-set principle, we characterize constructive faithfulness for Cantor series expansions using a log limit condition of the terms appearing in the series. This generalizes the invariance result of constructive dimension under base $b$ representations to all Cantor series expansions that obey this log limit condition. Moreover, it implies that for any Cantor series expansion that does not obey the log limit condition, there are sequences whose Cantor series dimension is different from its constructive dimension.
The recent works of J. Lutz, N. Lutz, Stull, Mayordomo and others study the “point-to-set principle” of how constructive Hausdorff dimension of points may be used to compute the classical Hausdorff dimension of arbitrary sets. In addition to the generalization of this point-to-set principle to $\Phi$-systems, our final result may be viewed as a new point-to-set phenomenon for the notion of “faithfulness”: here, equality of the constructive Cantor series dimensions and constructive dimensions of every point yield equality for the classical Cantor series and Hausdorff dimensions of every set, and conversely.
2 Preliminaries
2.1 Notation
We use $\Sigma$ to denote the binary alphabet $\{0,1\}$, $\Sigma^{*}$ represents the set of finite binary strings, and $\Sigma^{\infty}$ represents the set of infinite binary sequences. We use $\lvert x\rvert$ to denote the length of a finite string $x\in\Sigma^{*}$. For an infinite sequence $X=X_{0}X_{1}X_{2}\dots$, we use $X\upharpoonleft n$ to denote the finite string consisting of the first $n$ symbols of $X$. When $n\geq m$ we also use the notation $X[m,n]$ to denote the substring $X_{m}X_{m+1}\dots X_{n}$ of $X\in\Sigma^{\infty}$. We call two sets $U$ and $V$ to be incomparable if $U\not\subseteq V$ and $V\not\subseteq U$. For a set $U\subseteq\mathbb{R}$, we denote $|U|$ to denote the diameter of $U$, that is $|U|=\sup_{x,y\in U}d(x,y)$, where $d$ is the Euclidean metric. We use $\emptyset$ to denote the empty set, and we assume $|\emptyset|=0$. For any finite collection of sets $\{U\}$, we use #$\#(U)$ to denote the number of elements in $U$. Given infinite sequences $X_{1},\dots X_{n}$, we define the interleaved sequence $X_{1}\oplus X_{2}\oplus\dots X_{n}$ to be the interleaved sequence $X=X_{1}[0]X_{2}[0]\dots X_{n}[0]X_{1}[1]\dots X_{n}[1]\dots$. For some fixed $n\in\mathbb{N}$, we use $\mathbb{X}\subseteq\mathbb{R}^{n}$ to denote the metric space under consideration. We call a set of strings $\mathcal{P}\subset\Sigma^{*}$ to be prefix free if there are no two strings $\sigma,\tau\in\mathcal{P}$ such that $\sigma$ is a proper prefix of $\tau$. Given $n\in\mathbb{N}$ we use $[n]$ to denote $\{0,1,\dots n-1\}$. Kolmogorov Complexity represents the amount of information contained in a finite string. For more details on Kolmogorov Complexity, see [5], [11], [28], [31].
Definition 1.
The Kolmogorov complexity of $\sigma\in\Sigma^{*}$ is defined as $K(\sigma)=\min\limits_{\pi\in\Sigma^{*}}\left\{\lvert\pi\rvert\mid U(\pi)=\sigma\right\}$, where $U$ is a fixed universal prefix free Turing machine.
2.2 Hausdorff Dimension
The following definitions are originally given by Hausdorff [8]. We take the definitions from Falconer [7].
Definition 2 (Hausdorff [8]).
Given a set $\mathcal{F}\subseteq{\mathbb{X}}$, a collection of sets $\{U_{i}\}_{i\in\mathbb{N}}$ where for each $i\in\mathbb{N}$, $U_{i}\subseteq{\mathbb{X}}$ is called a $\delta$-cover of $\mathcal{F}$ if for all $i\in\mathbb{N}$, $|U_{i}|\leq\delta$ and $\mathcal{F}\subseteq\bigcup_{i\in\mathbb{N}}{U_{i}}$.
Definition 3 (Hausdorff [8]).
Given an $\mathcal{F}\subseteq{\mathbb{X}}$, for any $s>0$, define
$\mathcal{H}^{s}_{\delta}(\mathcal{F})=\inf\left\{\sum_{i}|U_{i}|^{s}:\{U_{i}\}_{i\in\mathbb{N}}\text{ is a }\delta\text{-cover of }\mathcal{F}\right\}.$
As $\delta$ decreases, the set of admissible $\delta$ covers decreases. Hence $\mathcal{H}^{s}_{\delta}(\mathcal{F})$ increases.
Definition 4 (Hausdorff [8]).
For $s\in(0,\infty)$, the s-dimensional Hausdorff outer measure of $\mathcal{F}$ is defined as:
$\mathcal{H}^{s}(\mathcal{F})=\lim\limits_{\delta\rightarrow 0}\;\mathcal{H}^{s}_{\delta}(\mathcal{F}).$
Observe that for any $t>s$, if $\mathcal{H}^{s}(\mathcal{F})<\infty$, then $\mathcal{H}^{t}(\mathcal{F})=0$ (see Section 2.2 in [7]).
Finally, we have the following definition of Hausdorff dimension.
Definition 5 (Hausdorff [8]).
For any $\mathcal{F}\subset{\mathbb{X}}$, the Hausdorff dimension of $\mathcal{F}$ is defined as:
${\mathrm{dim}}(\mathcal{F})=\inf\{s\geq 0:\mathcal{H}^{s}(\mathcal{F})=0\}.$
2.3 Constructive dimension
Lutz [12] defines the notion of effective (equivalently, constructive) dimension of an infinite binary sequence using the notion of lower semicomputable $s$-gales.
Definition 6 (Lutz [12]).
For $s\in[0,\infty)$, a binary $s$-gale is a function $d:\Sigma^{*}\to[0,\infty)$ such that $d(\lambda)<\infty$ and for all $w\in\Sigma^{*}$, $d(w)=2^{s}.\sum_{i\in\{0,1\}}d(wi).$
The success set of $d$ is $S^{\infty}(d)=\left\{X\in\Sigma^{\infty}\mid\limsup\limits_{n\to\infty}d(X\upharpoonleft n)=\infty\right\}$.
For $\mathcal{F}\subseteq[0,1]$, $\mathcal{G}(\mathcal{F})$ denotes the set of all $s\in[0,\infty)$ such that there exists a lower semicomputable (Definition 20) binary $s$-gale $d$ with $\mathcal{F}\subseteq S^{\infty}(d)$.
Definition 7 (Lutz [12]).
The constructive dimension or effective Hausdorff dimension of $\mathcal{F}\subseteq[0,1]$ is
$\mathrm{cdim}(\mathcal{F})=\inf\mathcal{G}(\mathcal{F}).$
The constructive dimension of a sequence $X\in\Sigma^{\infty}$ is $\mathrm{cdim}(X)=\mathrm{cdim}(\{X\})$.
Mayordomo [22] extends the result by Lutz [13] to give the following Kolmogorov complexity characterization of constructive dimension of infinite binary sequences.
Theorem 1 (Lutz [13], Mayordomo [22]).
For any $X\in\Sigma^{\infty}$,
$\mathrm{cdim}(X)=\liminf\limits_{n\to\infty}\frac{K(X\upharpoonleft n)}{n}.$
Mayordomo [23] later gave the following Kolmogorov complexity characterization of constructive dimension of points in $\mathbb{R}^{n}$.
Definition 8 (Mayordomo [23]).
For any $x\in\mathbb{R}^{n}$,
$\mathrm{cdim}(x)=\liminf\limits_{r\to\infty}\frac{K_{r}(x)}{r}.$
where $K_{r}(x)=\min\limits_{q\in\mathbb{Q}^{n}}\{K(q):|x-q|<2^{-r}\}$.
Constructive dimension also works in the Euclidean space. For a real $x=(x_{1},...x_{n})\in\mathbb{R}^{n}$, let binary expansions of the fractional parts of each of the coordinates of $x$ be $S_{1}\in\Sigma^{\infty},\dots,S_{n}\in\Sigma^{\infty}$ respectively. Then $\mathrm{cdim}(x)=n\cdot\mathrm{cdim}(X)$ where $X$ is the interleaved sequence $X=S_{1}\oplus S_{2}\dots\oplus S_{n}$ [17].
We now state some useful properties of constructive dimension. Lutz [12] shows that the constructive dimension of a set is always greater than or equal to its Hausdorff dimension.
Lemma 1 (Lutz [12]).
For any $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}(\mathcal{F})\leq\mathrm{cdim}(\mathcal{F})$.
Further, Lutz[12] also shows that the constructive dimension of a set is the supremum of the constructive dimensions of points in the set.
Lemma 2 (Lutz [12]).
For any $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\sup\limits_{x\in\mathcal{F}}\mathrm{cdim}(x).$
3 Hausdorff $\Phi$-dimension and Effective $\Phi$-dimension
Hausdorff dimension is defined using the notion of $s$-dimensional outer measures, where a cover is taken as the of union of a collection of covering sets $\{U_{i}\}_{i\in\mathbb{N}}$. Here a covering set $U_{i}$ can be any arbitrary subset of the space (see Section 2.2). We define the general notion of Hausdorff $\Phi$-dimension by restricting the class of admissible covers to $\Phi$-covers, which are the union of sets from a family of covering sets $\Phi$.
3.1 Family of covering sets
In this work, we consider a family of covering sets which satisfy the properties given below.
Definition 9 (Family of covering sets $\Phi$).
We consider the space ${\mathbb{X}}\subseteq\mathbb{R}^{\eta}$ where $\eta\in\mathbb{N}$. A countable family of sets $\Phi=\bigcup_{n\in\mathbb{Z}}\{U_{i}^{n}\}_{i\in\mathbb{N}}$, where for each $i\in\mathbb{N},n\in\mathbb{Z}$, $U_{i}^{n}\subseteq{\mathbb{X}}$, is called a family of covering sets if it satisfies the following properties:
  • Increasing Monotonicity: For every $n\in\mathbb{Z}$, $U\in\{U_{i}^{n}\}_{i\in\mathbb{N}}$ and $m\leq n$, there is a unique $V\in\{U_{i}^{m}\}_{i\in\mathbb{N}}$ such that $U\subseteq V$.
  • Fineness : Given any $\epsilon>0$, and $x\in{\mathbb{X}}$, there exists a $U\in\Phi$ such that $|U|<\epsilon$ and $x\in U$.
Note that the number of sets $\{U_{i}^{n}\}$ in some level $n\in\mathbb{Z}$ can also be finite and bounded by $m$. The definition still holds because in this case we take $U_{j}^{n}=\emptyset$ for $j>m$. Note that from Increasing monotonicity property, it follows that all elements $\{U_{i}^{n}\}$ in a particular level $n\in\mathbb{Z}$ are incomparable.
We now define the notion of a $\Phi$-cover of a set.
Definition 10 ($\Phi$-cover).
Let $\Phi=\bigcup_{n\in\mathbb{Z}}\{U_{i}^{n}\}_{i\in\mathbb{N}}$ be a family of covering sets. A $\Phi$ cover of a set $\mathcal{F}\subseteq{\mathbb{X}}$ is collection of sets $\{V_{j}\}_{j\in\mathbb{N}}\subseteq\Phi$ such that $\{V_{j}\}_{j\in\mathbb{N}}$ covers $\mathcal{F}$, that is $\mathcal{F}\subseteq\bigcup\limits_{j\in\mathbb{N}}V_{j}$
Note: Mayordomo [23] gives a definition of Nice covers of a metric space. They then give the definition of constructive dimension on a metric space with a nice cover. We note here that the notion of Family of covering sets is incomparable with the notion of Nice covers. Our definition does not require the $c$-cover property and the Decreasing monotonicity property of nice covers. Therefore, our notion includes the setting of continued fraction dimension, which is not captured by Nice covers. Also, the Fineness property required in our definition is not there in the definition of nice covers. The notion of Increasing monotonicity is present in both settings.
3.2 Hausdorff $\Phi$-dimension
Recall from Definition 10 that a $\Phi$-cover of $\mathcal{F}$ is a collection of sets from $\Phi$ that covers $\mathcal{F}$. We call this as a $\delta$-cover if the diameter of elements in the cover are less than $\delta$.
Definition 11.
Let $\Phi$ be a family of covering sets defined over ${\mathbb{X}}$. Given a set $\mathcal{F}\subseteq{\mathbb{X}}$, a $\Phi$-cover $\{U_{i}\}_{i\in\mathbb{N}}$ of $\mathcal{F}$ is called a $\delta$-cover of $\mathcal{F}$ using $\Phi$ if for all $i\in\mathbb{N}$, $|U_{i}|\leq\delta$.
Definition 12.
Given an $\mathcal{F}\subseteq{\mathbb{X}}$, for any $s>0$, we define
$\mathcal{H}^{s}_{\delta}(\mathcal{F},\Phi)=\inf\left\{\sum_{i}|U_{i}|^{s}:\{U_{i}\}_{i\in\mathbb{N}}\text{ is a }\delta\text{-cover of }\mathcal{F}\text{ using }\Phi\right\}.$
From the fineness property given in Definition 9, it follows that for any $\mathcal{F}\subseteq{\mathbb{X}}$, and $\delta>0$, $\delta$-covers of $\mathcal{F}$ using $\Phi$ always exist.
As $\delta$ decreases, the set of admissible $\delta$-covers using $\Phi$ decreases. Hence $\mathcal{H}^{s}_{\delta}(\mathcal{F},\Phi)$ increases.
Definition 13.
For $s\in(0,\infty)$, define the s-dimensional $\Phi$ outer measure of $\mathcal{F}$ as:
$\mathcal{H}^{s}(\mathcal{F},\Phi)=\lim\limits_{\delta\rightarrow 0}\;\mathcal{H}^{s}_{\delta}(\mathcal{F},\Phi).$
Observe that as with the case of classical Hausdorff dimension, for any $t>s$, if $\mathcal{H}^{s}(\mathcal{F},\Phi)<\infty$, then $\mathcal{H}^{t}(\mathcal{F},\Phi)=0$ (see Section 2.2 in [7]).
Finally, we have the following definition of Hausdorff $\Phi$-dimension.
Definition 14.
For any $\mathcal{F}\subset{\mathbb{X}}$, the Hausdorff $\Phi$-dimension of $\mathcal{F}$ is defined as:
${\mathrm{dim}}_{\Phi}(\mathcal{F})=\inf\{s\geq 0:\mathcal{H}^{s}(\mathcal{F},\Phi)=0\}.$
3.3 Effective $\Phi$-dimension
We first formulate the notion of a $\Phi$-$s$-supergale. A $\Phi$-$s$-supergale can be seen as a gambling strategy where the bets are placed on the covering sets from $\Phi$. The definitions in this subsection are adaptations from Mayordomo [23].
Definition 15 (Mayordomo [23]).
Let $\Phi=\bigcup_{n\in\mathbb{Z}}\{U_{i}^{n}\}_{i\in\mathbb{N}}$ be a family of covering sets from Definition 9. For $s\in[0,\infty)$, a $\Phi$-$s$-supergale is a function $d:\Phi\to[0,\infty)$ such that:
  • $\sum\limits_{U\in\{U_{i}^{0}\}_{i\in\mathbb{N}}}d(U)|U|^{s}<\infty$ and
  • For all $n\in\mathbb{N}$ and all $U\in\{U_{i}^{n}\}_{i\in\mathbb{N}}$, the following condition holds:
    $d(U).|U|^{s}\geq\sum_{V\in\{U_{i}^{n+1}\}_{i\in\mathbb{N}},V\subseteq U}d(V)|V|^{s}\;.$
The following is the generalization of Kraft inequality for $s$-supergales from Mayordomo [23].
Lemma 3 (Generalisation of Kraft inequality [23]).
Let $d$ be a $\Phi$-s-supergale. Then for every $\mathcal{E}\subseteq\Phi$ such that the sets in $\mathcal{E}$ are incomparable, we have that
$\sum\limits_{V\in\mathcal{E}}d(V)|V|^{s}\leq\sum\limits_{U\in\{U_{i}^{0}\}_{i\in\mathbb{N}}}d(U)|U|^{s}.$
Proof.
Let $\sum\limits_{U\in\{U_{i}^{0}\}_{i\in\mathbb{N}}}d(U)|U|^{s}=c$ for some $c\in\mathbb{R}$. It suffices to show that the Lemma holds when sets in $\mathcal{E}$ are subsets of level sets till a finite level $N$. That is, there exists an $N\in\mathbb{N}$ such that for all $U\in\mathcal{E}$, $U\in\{U_{i}^{m}\}$ for some $m\leq N$. In this case, we show that the lemma follows from induction. The base case when $N=0$ is immediate. Now assume the lemma holds for $N-1$, consider $\mathcal{E}_{N}=\{V\in\mathcal{E}\;|\;V\in\{U_{i}^{N}\}\}$ and $\mathcal{E}_{<N}=\mathcal{E}\setminus\mathcal{E}_{N}$. From the increasing monotonicity property in Definition 9, we have that for all $V\in\mathcal{E}_{N}$, $V\subseteq W$ for a unique $W\in\{U_{i}^{N-1}\}$. Let $\mathcal{E}_{N-1}^{\prime}$ be the collection of such $W$’s. We have that elements in $\mathcal{E}_{<N}\cup\mathcal{E}_{N-1}^{\prime}$ are incomparable and so $\sum\limits_{V\in\mathcal{E}_{<N}\cup\mathcal{E}_{N-1}^{\prime}}d(V)|V|^{s}\leq c$. From the gale condition in Definition 15, we have that for all $W\in\mathcal{E}_{N-1}^{\prime}$, $d(W).|W|^{s}\geq\sum_{V\in\{U_{i}^{N}\},V\subseteq U}d(V)|V|^{s}\;$. It follows that $\sum\limits_{U\in\mathcal{E}}d(U)|U|^{s}\leq c$.
Definition 16 (Mayordomo [23]).
Given $x\in{\mathbb{X}}$, a $\Phi$-representation of $x$ is a sequence $(U_{n})_{n\in\mathbb{Z}}$ such that for each $n\in\mathbb{Z}$, $U_{n}\in\{U_{i}^{n}\}_{i\in\mathbb{N}}$ and $x\in\cap_{n}U_{n}$.
Note that the same $x$ can have multiple $\Phi$-representations. Given $x\in{\mathbb{X}}$, let $\mathcal{R}(x)$ be the set of $\Phi$-representations of $x$.
Definition 17 (Mayordomo [23]).
A $\Phi$-s-supergale $d$ succeeds on $x\in{\mathbb{X}}$ if there is a $(U_{n})_{n\in\mathbb{Z}}\in\mathcal{R}(x)$ such that $\limsup\limits_{n\rightarrow\infty}d(U_{n})=\infty.$
Equivalently, a $\Phi$-$s$-supergale $d$ succeeds on a point $x\in{\mathbb{X}}$ iff for every $k\in\mathbb{N}$, there exists a $U\in\Phi$ such that $x\in U$ and $d(U)>2^{k}$.
Definition 18.
The success set of $d$ is $S^{\infty}(d)=\left\{x\in{\mathbb{X}}\mid d\text{ succeeds on x}\right\}.$
To define constructive $\Phi$-dimension, we require some additional computability restrictions over $\Phi$. It is an adaptation of the definition from [23].
Definition 19 (Family of computable covering sets $\Phi$).
We consider the space ${\mathbb{X}}\subseteq\mathbb{R}^{\eta}$ where $\eta\in\mathbb{N}$ and a family of covering sets $\Phi=\bigcup_{n\in\mathbb{Z}}\{U_{i}^{n}\}_{i\in\mathbb{N}}$ from Definition 9. We call $\Phi$ to be a family of computable covering sets if it satisfies the following additional properties:
  • Computable diameter: For every $n\in\mathbb{Z}$ and $i\in\mathbb{N}$, $|U_{i}^{n}|$ is uniformly computable.
  • Computable subsets: For every $n\in\mathbb{Z}$, and $i\in\mathbb{N}$, the set $\{j\in\mathbb{N}:U_{j}^{n+1}\subseteq U_{i}^{n}\}$ is uniformly computable.
In definition 19, when we say $|U_{i}^{n}|$ is uniformly computable, we mean that there is a turing machine that on input $n,i,r$ outputs a $q\in Q$ such that $||U_{i}^{n}|-q|<2^{r}$. The set $\{j\in\mathbb{N}:U_{j}^{n+1}\subseteq U_{i}^{n}\}$ is uniformly computable if there is a turing machine which on input $i,j,n$ decides if $U_{j}^{n+1}\subseteq U_{i}^{n}$.
We use constructive $\Phi$-$s$-gales to define the notion of constructive $\Phi$-dimension. For a $\Phi$-$s$-gale $d$ to be constructive, we require the gale function $d$ to be lower semicomputable. Note that a lower semicomputable supergale actually takes as input $(i,n)$ where $i\in\mathbb{N},n\in\mathbb{Z}$ to place bets on $U_{i}^{n}$. We omit this technicality in this paper and keep the domain of the gale as $\Phi$ for the sake of simplicity.
Definition 20.
A function $d:\Phi\longrightarrow[0,\infty)$ is called lower semicomputable if there exists a total computable function $\hat{d}:\Phi\times\mathbb{N}\longrightarrow\mathbb{Q}\cap[0,\infty)$ such that the following two conditions hold.
  • Monotonicity : For all $U\in\Phi$ and for all $n\in\mathbb{N}$, we have $\hat{d}(U,n)\leq\hat{d}(U,n+1)\leq d(U)$.
  • Convergence : For all $U\in\Phi$, $\lim\limits_{n\to\infty}\hat{d}(U,n)=d(U)$.
For $\mathcal{F}\subseteq{\mathbb{X}}$, let $\mathcal{G}_{\Phi}(\mathcal{F})$ denote the set of all $s\in[0,\infty)$ such that there exists a lower semicomputable $\Phi$-$s$-supergale $d$ with $\mathcal{F}\subseteq S^{\infty}(d)$.
Definition 21.
The constructive $\Phi$-dimension of $\mathcal{F}\subseteq{\mathbb{X}}$ is
$\mathrm{cdim}_{\Phi}(\mathcal{F})=\inf\mathcal{G}_{\Phi}(\mathcal{F}).$
The constructive $\Phi$ dimension of a point $x\in{\mathbb{X}}$ is defined by $\mathrm{cdim}_{\Phi}(\{x\})$, the constructive $\Phi$-dimension of the singleton set containing $x$.
This definition can easily be relativized with respect to an oracle $A\subseteq\mathbb{N}$ by giving the $s$-supergale an additional oracle access to set $A\subseteq\mathbb{N}$. We denote this using $\mathrm{cdim}_{\Phi}^{A}(\mathcal{F})$.
We now show that the constructive $\Phi$-dimension of a set is the supremum of constructive $\Phi$-dimensions of points in the set. The proof is a straightforward adaptation of proof of Theorem 2 by Lutz [12].
Theorem 2.
For any family of computable covering sets $\Phi$ defined over the space ${\mathbb{X}}$, for any $\mathcal{F}\subseteq{\mathbb{X}}$, we have
$\mathrm{cdim}_{\Phi}(\mathcal{F})=\sup\limits_{x\in\mathcal{F}}\mathrm{cdim}_{\Phi}(x).$
Proof.
Let $s$ be a rational such that $s>\mathrm{cdim}_{\Phi}(\mathcal{F})$. From the definition of $\mathrm{cdim}_{\Phi}(\mathcal{F})$ in Definition 9, there exists a lower semicomputable $\Phi$-$s$-supergale that succeeds on every $x\in\mathcal{F}$. So, for every $x\in\mathcal{F}$ and every rational $s>\mathrm{cdim}_{\Phi}(\mathcal{F})$, we have that $s\geq\mathrm{cdim}_{\Phi}(x)$. It follows that for every $x\in\mathcal{F}$, $\mathrm{cdim}_{\Phi}(\mathcal{F})\geq\mathrm{cdim}_{\Phi}(x)$ Therefore, $\mathrm{cdim}_{\Phi}(\mathcal{F})\geq\sup_{x\in\mathcal{F}}\mathrm{cdim}_{\Phi}(x)$.
To prove the opposite inequality consider any rational $s>\sup_{x\in\mathcal{F}}\mathrm{cdim}_{\Phi}(x)$. We show that $\mathrm{cdim}_{\Phi}(\mathcal{F})\leq s$. This implies the required inequality. Towards this we show that there exists a $\Phi$-$s$-supergale $d$ such that $d$ succeeds on every point $x\in\mathcal{F}$. This is done by enumerating all lower semicomputable $\Phi$-$s$-supergales and considering their convex combination. Since $s$ is rational, this enumeration is done in the same manner as in Theorem 6.3.7 in [6]. Having produced an enumeration $\{d_{i}\}_{i\in\mathbb{N}}$ of all $s$-supergales, consider $d$ defined as $d(U)=\sum_{i\in\mathbb{N}}d_{i}(U)2^{-i}$. It is straightforward to verify that $d$ is a lower semicomputable $\Phi$-$s$-supergale. Also, $S^{\infty}(d)\supseteq\cup_{i\in\mathbb{N}}S^{\infty}(d_{i})$. Since for every point $x\in\mathcal{F}$ there exists a corresponding lower semicomputable $\Phi$-$s$-supergale $d_{i}$ that succeeds on $x$, $d$ succeeds on every point in $\mathcal{F}$. ∎
4 Point-to-set principle for $\Phi$-dimension
Let $\Phi=\bigcup_{n\in\mathbb{Z}}\{U_{i}^{n}\}_{i\in\mathbb{N}}$ be a family of computable covering sets from Definition 19. In this work, we introduce the Point-to-Set principle for $\Phi$-dimension. We show that the Hausdorff $\Phi$-dimension of any set $\mathcal{F}\subseteq{\mathbb{X}}$ is equal to the relative constructive $\Phi$-dimensions of elements in the set, relative to a minimizing oracle $A$.
We first show that Hausdorff $\Phi$-dimension of any set is greater than or equal to the supremum of the relative constructive dimensions of elements in the set, over a minimizing oracle $A$. The main idea of the proof is that $A$ encodes the indices of $s$-dimensional outer covers of $\mathcal{F}$ for all $s\in\mathbb{Q}$ having $s>{\mathrm{dim}}_{\Phi}(\mathcal{F})$.
Lemma 4.
${\mathrm{dim}}_{\Phi}(\mathcal{F})\geq\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(x)$
Proof.
By definition of $\Phi$-dimension, we have that for any $s>{\mathrm{dim}}_{\Phi}(\mathcal{F})$ and $r\in\mathbb{N}$, there exists a sequence of $\Phi$-covers $\{{\mathcal{W}}_{i}^{r,s}\}_{i\in\mathbb{N}}$ of $\mathcal{F}$ such that $\sum_{i\in\mathbb{N}}|{\mathcal{W}}_{i}^{r,s}|^{s}\leq 2^{-r}$.
Consider a sequence $A\subseteq\mathbb{N}$ that encodes the indices of the cylinders $\{{\mathcal{W}}_{i}^{r,s}\}_{i\in\mathbb{N}}$ for each $i\in\mathbb{N}$, $r\in\mathbb{N}$ and for all $s\in\mathbb{Q}\;\cap\;({\mathrm{dim}}_{\Phi}(\mathcal{F}),\infty)$. We show that for any $s\in({\mathrm{dim}}_{\Phi}(\mathcal{F}),\infty)\cap\mathbb{Q}$, and for all $x\in\mathcal{F}$, $\mathrm{cdim}^{A}_{\Phi}(x)\leq s$. Given $r\in\mathbb{N}$, and $s\in({\mathrm{dim}}_{\Phi}(\mathcal{F}),\infty)\cap\mathbb{Q}$, consider the following $\Phi$-$s$-supergale $d_{r}:\Phi\to[0,\infty)$
$d_{r}(U)=\frac{1}{|U|^{s}}{\sum_{V\in\{{\mathcal{W}}_{i}^{r}\}\;;\;V\subseteq U}|V|^{s}}\;.$
It is routine to verify that $d_{r}$ is a $\Phi$-$s$-supergale. Note that $\sum_{U\in\{U_{i}^{0}\}}d(U)|U|^{s}=\sum_{V\in\{{\mathcal{W}}_{i}^{r}\}}\mu^{s}(V)\leq 2^{-r}$. Since $|V|$ is computable, $d_{r}$ is lower semicomputable when given an oracle access to $A$. Note that for all $U\in\{{\mathcal{W}}_{i}^{r}\}$, $d_{r}(U)\geq 1$. Finally define
$d(U)=\sum_{r=0}^{\infty}2^{r}d_{2r}(U).$
Now, for all $x\in\mathcal{F}$ and $r\in\mathbb{N}$, there exists a $U\in\{{\mathcal{W}}_{i}^{2r,s}\}$ such that $x\in U$. For that $U$, $d(U)\geq 2^{r}d_{2r}(U)=2^{r}$. Therefore $\mathcal{F}\subseteq S^{\infty}(d)$ and so $\mathrm{cdim}^{A}_{\Phi}(\mathcal{F})\leq s$. ∎
We now show that the Hausdorff dimension of a set is less than or equal to the supremum of the relative constructive dimensions of elements in the set over any oracle $A$.
Lemma 5.
${\mathrm{dim}}_{\Phi}(\mathcal{F})\leq\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(x)$
Proof.
Given any $A\subseteq\mathbb{N}$, let $s>\sup_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(x)$. From Theorem 2, it follows that there exists a lower semicomputable $\Phi$-$s$-gale, say $d$, that given an oracle access to $A$, succeeds on all $x\in\mathcal{F}$.
Given $r\in\mathbb{N}$ and $10<\delta<$ 1, define
$\mathcal{V}_{r}^{\delta}=\big{\{}U\in\Phi\;|\;d(U)\geq\frac{2^{r}}{\delta^{s}}\big{\}}.$
Let ${\mathcal{W}}_{r}^{\delta}$ be the subset of $\mathcal{V}_{r}^{\delta}$ such that elements in ${\mathcal{W}}_{r}^{\delta}$ are incomparable and $\bigcup_{U\in\mathcal{V}_{r}^{\delta}}=\bigcup_{U\in{\mathcal{W}}_{r}^{\delta}}$. Since $\mathcal{F}\subseteq S^{\infty}[d]$, it follows for all $r\in\mathbb{N}$ and $\delta>0$, $\mathcal{F}\subseteq\bigcup\limits_{U\in{\mathcal{W}}_{r}^{\delta}}U$. From Lemma 3, it follows that for all $U\in{\mathcal{W}}_{r}^{\delta}$, $|U|<\delta$.
From Lemma 3, it also follows that for some $c\in\mathbb{N}$,
$\sum\limits_{U\in{\mathcal{W}}_{r}^{\delta}}d(U)|U|^{s}\leq c$
We have that for all $U\in{\mathcal{W}}_{r}^{\delta}$, $d(U)\geq 2^{r}$. Therefore, it follows that
$\sum\limits_{U\in{\mathcal{W}}_{r}^{\delta}}|U|^{s}\leq c.2^{-r}.$
Hence for all $\delta>0$, and $r\in\mathbb{N}$, $H^{s}_{\delta}(\mathcal{F},\Phi)\leq c.2^{-r}$. Therefore we have that $H^{s}_{\delta}(\mathcal{F},\Phi)=0$. So it follows that ${\mathrm{dim}}_{\Phi}(\mathcal{F})\leq s$. ∎
Combining Lemma 4 and Lemma 5, we have the point-to-set principle for $\Phi$-dimension.
Theorem 3.
For a family of computable covering sets $\Phi$ over the space ${\mathbb{X}}$, for all $\mathcal{F}\subseteq{\mathbb{X}}$,
${\mathrm{dim}}_{\Phi}(\mathcal{F})=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(x).$
4.1 Point to Set Principle for constructive dimension
Definition 22 (Dyadic Family of covers).
Consider the space ${\mathbb{X}}=\mathbb{R}^{n}$. The dyadic family of covers is the set of coverings $\Phi_{B}=\bigcup_{r\in\mathbb{N}}\{[\frac{m_{1}}{2^{r}},\frac{m_{1}+1}{2^{r}}]\times\dots\times[\frac{m_{n}}{2^{r}},\frac{m_{n}+1}{2^{r}}]\}_{m_{1},m_{2}\dots,m_{n}\in[2^{r}]}$.
It is straightforward to verify that $\Phi_{B}$ is a family of computable covering sets from Definition 19. Besicovitch [3] gave the following $\Phi$-dimension characterization of Hausdorff dimension.
Lemma 6 (Besicovitch [3]).
For all $\mathcal{F}\subseteq\mathbb{R}^{n}$, we have ${\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{\Phi_{B}}(\mathcal{F}).$
Similarly, we have the following $\Phi$-dimension characterization of Constructive dimension.
Lemma 7 (Lutz and Mayordomo [17]).
For all $\mathcal{F}\subseteq\mathbb{R}^{n}$, we have $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi_{B}}(\mathcal{F}).$
From Theorem 3 for $\Phi_{B}$ and using Lemma 6 and 7, we have the following point-to-set principle from [14] relating Hausdorff and Constructive dimensions.
Corollary 1 (J.Lutz and N.Lutz [14]).
For all $\mathcal{F}\subseteq{\mathbb{X}}$,
${\mathrm{dim}}(\mathcal{F})=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}(x).$
4.2 Point to Set Principle for Continued Fraction dimension
The sequence $Y=[a_{1},a_{2},\dots]$ where each $a_{i}\in\mathbb{N}$ is the continued fraction expansion of the number $y=\cfrac{1}{a_{1}+\cfrac{1}{a_{2}+\cdots}}$. Given $u=[a_{1},a_{2}\dots a_{n}]\in\mathbb{N}^{*}$, the cylinder set of $u$, $C_{u}$ is defined as $C_{u}=[[a_{1},a_{2},\dots a_{n}],[a_{1},a_{2},\dots a_{n}+1]]$ when $n$ is even and $C_{u}=[[a_{1},a_{2},\dots a_{n}+1],[a_{1},a_{2},\dots a_{n}]]$ when $n$ is odd.
The notion of constructive continued fraction dimension was introduced by Nandakumar and Vishnoi [27] using continued fraction $s$-gales. Akhil, Nandakumar and Vishnoi [25] showed that this notion is different from that of constructive dimension.
Consider $\Phi_{CF}$ to be the set of covers generated by the continued fraction cylinders, that is $\Phi_{CF}=\bigcup_{n\in\mathbb{Z}}\{C_{[a_{1},a_{2},\dots a_{n}]}\}_{a_{1}\dots a_{n}\in\mathbb{N}}$. It is routine to verify that this is a family of computable covering sets from Definition 19. From Theorem 3, we therefore have the following point-to-set principle for Continued fraction dimension.
Corollary 2.
For all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}_{\text{CF}}(\mathcal{F})=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\text{CF}}(x)$.
4.3 Effective $\Phi$-dimension using Kolmogorov Complexity
We give an equivalent formulation of constructive $\Phi$-dimension of a point using Kolmogorov complexity. For this, we require some additional properties for the space $\Phi$.
Definition 23 (Family of finitely intersecting computable covering sets $\Phi$).
We consider the space ${\mathbb{X}}\subseteq\mathbb{R}^{\eta}$ where $\eta\in\mathbb{N}$ and a family of computable covering sets $\Phi=\bigcup_{n\in\mathbb{Z}}\{U_{i}^{n}\}_{i\in\mathbb{N}}$ from Definition 19. We say that $\Phi$ is a family of finitely intersecting computable covering sets if it satisfies the following additional properties:
  • Density of Rational points: For each $U\in\Phi$, there exists a $q\in\mathbb{Q}^{n}$ such that $q\in U$.
  • Finite intersection: There exists a constant $c\in\mathbb{N}$ such that for any collection $\{U_{i}\}\subseteq\Phi$ satisfying (1) $U_{i}\not\subseteq U_{j}$for all $i\neq j$, and (2) $\bigcap_{i}U_{i}\neq\emptyset$,
    we have #($\#(\{U_{i}\})\leq c$.
  • Membership test: There is a computable function that given $i\in\mathbb{N}$ and $n\in\mathbb{Z}$ and a $q\in\mathbb{Q}^{n}$ checks if $q\in U_{i}^{n}$.
The Finite intersection property states that the cardinality of any collection of incomparable sets from $\Phi$ having non empty intersection is bounded by a constant.
Given family of finitely intersecting computable covering sets $\Phi$ over a space ${\mathbb{X}}$, and an $r\in\mathbb{N}$, we define the notion of Kolmogorov Complexity of a point $X$ at precision $r$ with respect to $\Phi$. We denote this using $K_{r}(X,\Phi)$.
Definition 24.
Given an $r\in\mathbb{N}$, and $x\in{\mathbb{X}}$, define
$K_{r}(x,\Phi)=\min_{U\in\Phi}\{K(U)\;|\;x\in U\text{ and }|U|<2^{-r}\}.$
where for $U\in\Phi$, $K(U)$ is defined as $K(U)=\min\{K(q)\;|\;q\in U\cap\mathbb{Q}^{n}\}.$
We now go on to prove a Kolmogorov Complexity characterization of constructive $\Phi$-dimension for a family of finitely intersecting computable covering sets $\Phi$.
Lemma 8.
Given a family of finitely intersecting computable covering sets $\Phi$, over a space ${\mathbb{X}}$. For any $x\in{\mathbb{X}}$,
$\mathrm{cdim}_{\Phi}(x)\leq\liminf_{r\rightarrow\infty}\frac{K_{r}(x,\Phi)}{r}$
Proof.
Given any $q\in{\mathbb{X}}$ and $k\in\mathbb{N}$, define $B_{k}(q)$ to the collection of maximal sets in $\Phi$ having size less than or equal to $2^{-k}$ containing $q$.
$B_{k}(q)=\{U\in\Phi\;\lvert\;|U|<2^{-k}\text{ and }q\in U,\\ \text{ and for all }V\in\Phi\text{ such that }V\neq U,|V|<2^{-k}\text{ and }q\in U,\text{ we have }U\not\subseteq V\}.$
Note that for any $q\in{\mathbb{X}}$, $B_{k}(q)\neq\emptyset$ due to the fineness condition in Definition 9. From the finite intersection property in definition 23, it follows that #$\#(B_{k}(q))\leq c$. From the membership test condition in Definition 23 and the Increasing Monotonicity property in Definition 9, we have that $B_{k}(q)$ is computable given $k$ and $q$.
For any $k\in\mathbb{N}$ and $s^{\prime\prime}\in\mathbb{Q}$, define the set $Q_{k}^{s^{\prime\prime}}=\{q\in\mathbb{Q}^{n}:K(q)<ks^{\prime\prime}\}$. We have that $Q_{k}^{s^{\prime\prime}}$ is computably enumerable.
For any $U\in\Phi$ and $s,s^{\prime},s^{\prime\prime}\in\mathbb{Q}$ such that $s>s^{\prime}>s^{\prime\prime}$, define
$d_{k}(U)=\frac{1}{|U|^{s}}{\sum_{q\in Q_{k}^{s^{\prime\prime}}}\sum_{W\in B_{k}(q);\;W\subseteq U}|W|^{s^{\prime}}}\;.$
It is routine to verify that $d_{k}:\Phi\to[0,\infty)$ is a lower semicomputable $\Phi$-$s$-supergale. Also, we have that $\sum_{U\in\{U_{i}^{0}\}}d_{k}(U)|U|^{s}\leq\sum_{q\in Q_{k}^{s^{\prime\prime}}}\sum_{W\in B_{k}(q)}|W|^{s^{\prime}}.$ Since there are fewer than $2^{ks^{\prime\prime}+1}$ $q$’s such that $K(q)\leq ks^{\prime\prime}$, we have $\sum_{U\in\{U_{i}^{0}\}}d_{k}(U)|U|^{s}\leq 2c.2^{k(s^{\prime\prime}-s^{\prime})}$.
Define
$d(U)=\sum_{k\in\mathbb{N}}d_{k}(U).$
We have that $\sum_{U\in\{U_{i}^{0}\}}d(U)|U|^{s}\leq 2c\sum_{k\in\mathbb{N}}2^{k(s^{\prime\prime}-s^{\prime})}<\infty$. Therefore, $d$ is a lower semicomputable $\Phi$-$s$-supergale.
Now given $x\in{\mathbb{X}}$, let $s,s^{\prime},s^{\prime\prime}\in\mathbb{Q}$ such that $s>s^{\prime}>s^{\prime\prime}>\liminf_{r\rightarrow\infty}\frac{K_{r}(x,\Phi)}{r}$. We have that for every $r\in\mathbb{N}$, there exists a $k\geq r$, and a $U\in\Phi$ having $|U|<2^{-k}$ and a $q\in\mathbb{Q}^{n}$ such that $K(q)<ks^{\prime\prime}$ and $x,q\in U$. We also have that $U\subseteq V$ for some $V\in B_{k}(q)$. Since $d(V)\geq|V|^{s^{\prime}-s}>2^{k(s-s^{\prime})}$, we have that the lower semicomputable $\Phi$-$s$-supergale $d$ succeeds on $x$. Therefore $\mathrm{cdim}_{\Phi}(x)\leq s$.
Lemma 9.
Given a family of finitely intersecting computable covering sets $\Phi$, over a space ${\mathbb{X}}$. For any $x\in{\mathbb{X}}$,
$\mathrm{cdim}_{\Phi}(x)\geq\liminf_{r\rightarrow\infty}\frac{K_{r}(x,\Phi)}{r}$
Proof.
Take any $s\in\mathbb{Q}$ such that $s>\mathrm{cdim}_{\Phi}(x)$. From the definition, it follows that there exists a lower semicomputable $\Phi$-$s$-supergale, say $d$, which succeeds on $x$.
Given $k,r\in\mathbb{N}$ , define
$\mathcal{V}_{k}^{r}=\big{\{}U\in\Phi\;|\;d(U)\geq{2^{ks}}\text{ and }|U|\in[2^{-r-1},2^{-r})\big{\}}.$
The set $\mathcal{V}_{k}^{r}$ is computably enumerable as $d$ is lower semicomputable and $|U|$ is computable from Definition 19. Let ${\mathcal{W}}_{k}^{r}$ be the set obtained by enumerating the sets in $\mathcal{V}_{k}$ and discarding the sets which are comparable to the ones enumerated so far.
From Lemma 3, it follows that #$\#({\mathcal{W}}_{k}^{r})\leq 2^{-ks+rs+s+O(1)}$. Also, from Lemma 3, it follows that for all $U\in{\mathcal{V}}_{k}$, $|U|\leq 2^{-k+O(1)}$. So if $\mathcal{V}_{k}^{r}\neq\emptyset$, we have that $k\leq r$
It follows that any set $U\in{\mathcal{W}}_{k}^{r}$ can be identified given $k$ and $r$, each requiring at most $\log r$ bits, and an index into the set ${\mathcal{W}}_{k}^{r}$ which requires at most $rs+s+O(1)$ bits. From the membership test property in definition 23, it therefore follows that for all $U\in{\mathcal{W}}_{k}^{r}$ there exists a $q\in\mathbb{Q}^{n}$ such that $q\in U$ and $K(q)\leq rs+o(r)$.
Since $d$ succeeds on $x$, we have that for all $k\in\mathbb{N}$, there exists a $U\in\Phi$ such that $x\in U$ and $d(U)>2^{ks}$. Let $|U|\in[2^{-r-1},2^{-r})$ for some $r\in\mathbb{N}$. We have that $U\in{\mathcal{V}}_{k}^{r}$.
There are now three cases. Either $U\in{\mathcal{W}}_{k}^{r}$, or there exists a $W\supseteq U$ such that $W\in{\mathcal{W}}_{k}^{r}$ or a $V\subseteq U$ such that $V\in{\mathcal{W}}_{k}^{r}$. In the second case, we have that $x\in W$ and there exists a $q\in\mathbb{Q}^{n}$ such that $q\in W$ having $K(q)<rs+o(r)$. In the remaining two cases, we have that there exists a $q\in\mathbb{Q}^{n}$ such that $x\in U$, $q\in U$ and $K(q)<rs+o(r)$.
Therefore we have that for any $k\in\mathbb{N}$, there exists a $r\geq k$ such that for some $U\in\Phi$ having $x\in U$ and $|U|<2^{-r}$, we have there exists a $q\in\mathbb{Q}^{n}$ such that $q\in U$ and $K(q)<rs+o(r)$. So we have that that $\liminf_{r\rightarrow\infty}\frac{K_{r}(x,\Phi)}{r}\leq s$. ∎
For any family of finitely intersecting computable covering sets $\Phi$, we have the following Kolmogorov Complexity characterization of constructive $\Phi$-dimension.
Theorem 4.
Given a family of finitely intersecting computable covering sets $\Phi$, over a space ${\mathbb{X}}$. For any $x\in{\mathbb{X}}$,
$\mathrm{cdim}_{\Phi}(x)=\liminf_{r\rightarrow\infty}\frac{K_{r}(x,\Phi)}{r}$
5 Kolmogorov Complexity Construction
In this section we give a technical construction which is crucial in proving the results in section 6. Theorem 5 says that given an infinite sequence $X$ and an oracle $A$, for any oracle $B$, there exists a sequence $Y$ whose relativized Kolmogorov complexity (of prefixes) with respect to $B$ is similar to the relativized Kolmogorov complexity (of prefixes) of $X$ with respect to $A$.
Theorem 5.
For all $X\in\Sigma^{\infty}$ and $A\in\Sigma^{\infty}$, given a $B\in\Sigma^{\infty}$, there exists a $Y\in\Sigma^{\infty}$ such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$ and $\mathrm{cdim}^{B}(Y)=\mathrm{cdim}(Y)$.
Proof.
We construct $Y$ in a stage-wise manner. We start with the empty string $\sigma_{0}=\lambda$ and define finite strings $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}\dots$ inductively. Finally we set $Y=\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3}\dots$.
There exist a constant $c_{K}$ such that for sufficiently large n, there exists some $w\in\Sigma^{n}$ with, $K^{B}(w)\geq n-c_{K}$. Let $M$ be any sufficiently large natural number such that $\log M\geq c_{K}$. For any $m\geq 1$, we do the following in stage $m$. If $m=1$ let $s_{1}=0$ and $e_{1}=M^{2}$. Otherwise we set $s_{m}=(M+m-1)^{2}$ and $e_{m}=(M+m)^{2}$.
If $K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])\geq\log(M+m)$, then let $\sigma_{m}=w_{m}z_{m}$ where $w_{m}$ is any $K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])$-length string such that, $K^{B}(w_{m}\mid\sigma_{0}\sigma_{1}\dots\sigma_{m-1})\geq K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])-c_{K}$ and $z_{m}$ is the string consisting of $e_{m}-s_{m}-K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])$ number of zeroes. If $K^{A}(X[s_{m}:e_{m}]\mid X[0:s_{m}-1])<\log(M+m)$ then we let $\sigma_{m}$ equal to the string consisting of $e_{m}-s_{m}$ number of zeroes. We define $Y=\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3}\dots$. We show that $Y$ satisfies the required properties.
To prove the first property, it is enough to show that
$\left\lvert K^{A}(X\upharpoonleft(M+m-1)^{2})-K^{B}(Y\upharpoonleft(M+m-1)^{2})\right\rvert\leq c(M+m-1)\log(M+m-1)^{2}.$
In the above $c=c_{K}+2+c_{sym}^{A}+c_{sym}^{B}$ where $c_{sym}^{A}$ and $c_{sym}^{B}$ are the constants from the asymptotic error term in the symmetry of information of Kolmogorov complexity with oracles $A$ and $B$ respectively (see [5]). The above statement implies that
$\left\lvert K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)\right\rvert\leq 4c\cdot\sqrt{n}\log n$
which proves the first part. The proof follows by induction on $n$. Consider the case when $m=1$. By the choice of $\sigma_{1}$ and the symmetry of information,
$K^{B}(Y\upharpoonleft M^{2})=K^{B}(w_{1})+K^{B}(z_{1}\mid w_{1})+O(\log M^{2})$
Now,
$\left\lvert K^{A}(X\upharpoonleft M^{2})-K^{B}(Y\upharpoonleft M^{2})\right\rvert\leq c_{K}+\log M^{2}+c_{sym}^{B}\log M^{2}\leq c\log M^{2}$
which proves the base case. Now, assume the statement for $n\geq 1$. Using the symmetry of information,
$K^{B}(Y\upharpoonleft(M+m)^{2})$ $=K^{B}(Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(z_{m}\mid Y\upharpoonleft(M+m-1)^{2},w_{m})+O(\log(M+m)^{2})$
Similarly,
$K^{A}(X\upharpoonleft(M+m)^{2})$ $=K^{A}(X\upharpoonleft(M+m-1)^{2})$
$+K^{A}(X[(M+m-1^{2}):(M+m)^{2}]\mid X\upharpoonleft(M+m-1)^{2})$
$+O(\log(M+m)^{2}).$
Hence, using the induction hypothesis,
$\left\lvert K^{A}(X\upharpoonleft(M+m-1)^{2})-K^{B}(Y\upharpoonleft(M+m-1)^{2})\right\rvert$
$\leq c(M+m-1)\log(M+m-1)^{2}+c_{K}$
$+2\log(M+m)^{2}+c_{sym}^{A}\log(M+m)^{2}+c_{sym}^{B}\log(M+m)^{2}$
$\leq c(M+m-1)\log(M+m)^{2}+(c_{K}+2+c_{sym}^{A}+c_{sym}^{B})\log(M+m)^{2}$
$=c(M+m)\log(M+m)^{2}.$
Now, we prove the second part. It is enough to show that,
$\left\lvert K(Y\upharpoonleft(M+m-1)^{2})-K^{B}(Y\upharpoonleft(M+m-1)^{2})\right\rvert\leq c^{\prime}(M+m-1)\log(M+m-1)^{2}.$
for every $m\geq 1$. Here $c^{\prime}=c_{K}+2+c_{sym}+c^{B}_{sym}$ where $c_{sym}$ is the constant in symmetry of information for non-relativized Kolmogorov complexity. This implies that
$\left\lvert K(Y\upharpoonleft n)-K^{B}(Y\upharpoonleft n)\right\rvert\leq 4c^{\prime}\cdot\sqrt{n}\log n$
which gives us the required conclusion. Consider the case when $m=1$. Now,
$K(Y\upharpoonleft M^{2})=K(w_{1})+K(z_{1}\mid w_{1})+O(\log M^{2}).$
By the choice of $w_{1}$, it is easy to verify that
$\left\lvert K(w_{1})-K^{B}(w_{1})\right\rvert\leq c_{K}.$
Therefore,
$\left\lvert K(Y\upharpoonleft M^{2})-K^{B}(Y\upharpoonleft M^{2})\right\rvert$ $\leq c_{K}+2\log M^{2}+c^{B}_{sym}\log M^{2}+c_{sym}\log M^{2}$
which proves the base case. Now, assume the statement for $m\geq 1$. We have,
$K^{B}(Y\upharpoonleft(M+m)^{2})$ $=K^{B}(Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})$
$+K^{B}(z_{m}\mid Y\upharpoonleft(M+m-1)^{2},w_{m})+O(\log(M+m)^{2}).$
Similarly,
$K(Y\upharpoonleft(M+m)^{2})$ $=K(Y\upharpoonleft(M+m-1)^{2})$
$+K(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})$
$+K(z_{m}\mid Y\upharpoonleft(M+m-1)^{2},w_{m})+O(\log(M+m)^{2})$
By the choice of $w_{m}$,
$\left\lvert K(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})-K^{B}(w_{m}\mid Y\upharpoonleft(M+m-1)^{2})\right\rvert\leq c_{K}$
Therefore using the induction hypothesis,
$\left\lvert K(Y\upharpoonleft(M+m)^{2})-K^{B}(Y\upharpoonleft(M+m)^{2})\right\rvert$
$\leq c^{\prime}(M+m-1)\log(M+m-1)^{2}+c_{K}$
$+2\log(M+m)^{2}+c_{sym}\log(M+m)^{2}+c_{sym}^{B}\log(M+m)^{2}$
$\leq c^{\prime}(M+m-1)\log(M+m)^{2}+(c_{K}+2+c_{sym}+c_{sym}^{B})\log(M+m)^{2}$
$=c^{\prime}(M+m)\log(M+m)^{2}.$
The proof of the second part is thus complete. ∎
6 Equivalence of Faithfulness of Cantor Coverings at Constructive and Hausdorff Levels
In this section, we show that when the class of covers $\Phi$ is generated by computable Cantor series expansions, the faithfulness at the Hausdorff and constructive levels are equivalent notions.
6.1 Faithfulness of Family of Coverings
We will first see the definition of Hausdorff dimension faithfulness. We then introduce the corresponding notion at the effective level, which we call constructive dimension faithfulness.
A family of covering sets $\Phi$ is said to be faithful with respect to Hausdorff dimension if the $\Phi$ dimension of every set in the space is the same as its Hausdorff dimension.
Definition 25.
A family of covering sets $\Phi$ over the space ${\mathbb{X}}$ is said to be faithful with respect to Hausdorff dimension if for all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}_{\Phi}(\mathcal{F})={\mathrm{dim}}(\mathcal{F})$.
We extend the definition to the constructive level as well. A family of computable covering sets $\Phi$ is defined to be faithful with respect to constructive dimension if the constructive $\Phi$ dimension of every set is the same as its constructive dimension.
Definition 26.
A family of computable covering sets $\Phi$ is said to be faithful with respect to constructive dimension if for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}_{\Phi}(\mathcal{F})=\mathrm{cdim}(\mathcal{F})$.
The following lemma follows from Theorem 2 and Lemma 2. It states that constructive dimension faithfulness can be equivalently stated in terms of preservation of constructive dimensions of points in the set.
Lemma 10.
A constructive family of covers $\Phi$ is faithful with respect to Constructive dimension if and only if for all $x\in{\mathbb{X}}$, $\mathrm{cdim}_{\Phi}(x)=\mathrm{cdim}(x)$.
The following lemma states that the $\Phi$-dimension of a set is always greater than or equal to its Hausdorff dimension. Similarly, the constructive $\Phi-$dimension of a set is always greater than or equal to its constructive dimension.
Lemma 11.
For any family of covering sets $\Phi$ over ${\mathbb{X}}$, for all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}_{\Phi}(\mathcal{F})\geq{\mathrm{dim}}(\mathcal{F})$.
Lemma 12.
For any family of finitely intersecting computable covering sets $\Phi$ over ${\mathbb{X}}$, for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}_{\Phi}(\mathcal{F})\geq\mathrm{cdim}(\mathcal{F})$.
Proof.
Consider any $x\in{\mathbb{X}}$ such that $s>\mathrm{cdim}_{\Phi}(x)$. We have that for infinitely many $r\in\mathbb{N}$, there exists a $U\in\Phi$ with $|U|<2^{-r}$ and a $q\in U$ having $K(q)<sr+o(r)$. Since $|U|<2^{-r}$, we have that $|x-q|<2^{-r}$. Since this happens for infinitely many $r$, we have that $\mathrm{cdim}(x)\leq s$. ∎
Therefore we have that $\Phi$ is not faithful for Hausdorff dimension if and only if there exists an $\mathcal{F}\subset{\mathbb{X}}$ such that ${\mathrm{dim}}_{\Phi}(\mathcal{F})>{\mathrm{dim}}(\mathcal{F})$. Similarly, $\Phi$ is not faithful for Constructive dimension if and only if there exists an $\mathcal{F}\subset{\mathbb{X}}$ such that $\mathrm{cdim}_{\Phi}(\mathcal{F})>\mathrm{cdim}(\mathcal{F})$.
6.2 Cantor coverings over unit interval
We consider the faithfulness of family of coverings generated by the computable Cantor series expansion [4]. We call such class of coverings as Cantor coverings.
Given a sequence $Q=\{n_{k}\}_{k\in\mathbb{N}}$ with $n_{k}\in\mathbb{N}\setminus\{1\}$, the expression
$x=\sum_{k=1}^{\infty}\frac{\alpha_{k}}{n_{1}.n_{2}\dots n_{k}}$
where $\alpha_{k}\in[n_{k}]$ is called the cantor series expansion of the real number $x\in[0,1]$.
Definition 27 (Cantor Coverings $\Phi_{Q}$).
The class of Cantor coverings $\Phi_{Q}$ over the space ${\mathbb{X}}=[0,1]$ generated by the Cantor series expansion $Q=\{n_{k}\}_{k\in\mathbb{N}}$ is the set of intervals
$\bigcup_{{k\in\mathbb{Z}}}\{[\frac{m}{n_{0}.n_{1}.n_{2}\dots n_{k}},\frac{m+1}{n_{0}.n_{1}.n_{2}\dots n_{k}}]\}_{m\in[n_{0}.n_{1}.n_{2}\dots n_{k}]}$
with $n_{0}$ taken as $1$.
Definition 28 (Computable Cantor Coverings).
The cantor series expansion $Q=\{n_{k}\}_{k\in\mathbb{N}}$ is said to be computable if there exists a machine that generates $n_{k}$ given $k$. We call the class of Cantor coverings $\Phi_{Q}$ generated by a computable Cantor series expansion $Q$ as a class of Computable Cantor Coverings over ${\mathbb{X}}=[0,1]$.
It is routine to verify that for any computable Cantor series expansion $Q=\{n_{k}\}_{k\in\mathbb{N}}$, the Cantor covering $\Phi_{Q}$ is a family of finitely intersecting computable covering sets from Definition 23. Therefore, from Theorem 3, we have the following Point-to-Set principle for Cantor covering dimension.
Corollary 3.
For all $\mathcal{F}\subseteq{\mathbb{X}}$ and for all computable Cantor coverings $\Phi_{Q}$,
${\mathrm{dim}}_{\Phi_{Q}}(\mathcal{F})=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi_{Q}}(x).$
6.3 Kolmogorov Complexity Characterization of Cantor Series Dimension
We first show a Kolmogorov complexity characterization of constructive $\Phi$-dimension for computable Cantor coverings.
Theorem 6.
For any $x\in{\mathbb{X}}$, and any computable Cantor coverings $\Phi_{Q}$ generated by $Q=\{n_{k}\}_{k\in\mathbb{N}}$,
$\mathrm{cdim}_{\Phi_{Q}}(x)=\liminf\limits_{k\to\infty}\frac{K(X\upharpoonleft m_{k})}{m_{k}}$
where $X$ is a binary expansion of $x$ and $m_{k}=\left\lfloor\log_{2}(n_{1}.n_{2}\dots n_{k})\right\rfloor.$
Proof.
For any $s>\mathrm{cdim}_{\Phi}(x)$, we have that for infinitely many $r\in\mathbb{N}$, there exists a $q\in\mathbb{Q}$ and a $U\in\Phi$ having $|U|<2^{-r}$ , such that $x,q\in U$ and $K(q)\leq rs+o(r)$. Let $U=[\frac{i}{n_{1}\dots n_{k}},\frac{i+1}{n_{1}\dots n_{k}}]$ for some $k\in\mathbb{N}$ where $i\in[n_{1}\dots n_{k}]$. From the definition of $m_{k}$, we have that $\frac{1}{2.2^{m_{k}}}<|U|\leq\frac{1}{2^{m_{k}}}$.
Consider a machine $M$ that takes in the input $\tau\kappa\delta$, each coming from a prefix free set of programs and produces $q,m_{k},e$ where $q\in\mathbb{Q},m_{k}\in\mathbb{N},e\in\{-1,0,1\}$. The machine outputs $\frac{j+e}{2^{m_{k}}}$ such that $q\in[\frac{j}{2^{m_{k}}},\frac{j+1}{2^{m_{k}}})$ where $j\in\{0,1\dots 2^{m_{k}}\}$.
Since $|U|\leq\frac{1}{2^{m_{k}}}$, we have that $x\in[\frac{j+e}{2^{m_{k}}},\frac{j+e+1}{2^{m_{k}}})$ for some $e\in\{-1,0,1\}$, and therefore $X\upharpoonleft m_{k}=\frac{j+e}{2^{m_{k}}}$. Since $|U|>\frac{1}{2.2^{m_{k}}}$ and $|U|<2^{-r}$, we have $m_{k}+1>r$. Since $K(q)<rs+o(r)$, we have $K(q)\leq s.m_{k}+o(m_{k})$ and so we have $K(X\upharpoonleft m_{k})\leq s.m_{k}+o(m_{k})$. Since this happens for infinitely many $k$, corresponding to each $r$, we have that $\liminf\limits_{k\to\infty}\frac{K(X\upharpoonleft m_{k})}{m_{k}}\leq s$.
Now consider any $s>\liminf\limits_{k\to\infty}\frac{K(X\upharpoonleft m_{k})}{m_{k}}$, we have that for infinitely many $k$, $K(X\upharpoonleft m_{k})\leq s.m_{k}+o(m_{k})$. Consider a machine that takes as the input $\tau\kappa\delta$, each coming from a prefix free set of programs and produces $q\in\mathbb{Q}$, $m_{k}\in\mathbb{N}$ and $e\in\{0,1,2\}$. The program then finds an $i$ such that $q=\frac{i}{2^{m_{k}}}$ and finally outputs $\frac{2i+e}{2.2^{m_{k}}}$.
Let $q=X\upharpoonleft m_{k}=\frac{i}{2^{m_{k}}}$. From the definition of $m_{k}$, we have that $\frac{1}{2.2^{m_{k}}}<|U|\leq\frac{1}{2^{m_{k}}}$ for some $U\in\Phi$ such that $x\in U$. As $|U|>\frac{1}{2.2^{m_{k}}}$, it follows that $p=\frac{2i+e}{2.2^{m_{k}}}\in U$ for some $e\in\{0,1,2\}$. Since $K(q)\leq s.m_{k}+o(m_{k})$, we have that $K(p)\leq s.m_{k}+o(m_{k})$. Therefore, $\mathrm{cdim}_{\Phi}(x)\leq s$. ∎
Theorem 1 ensures that when the Kolmogorov complexities of any two $X,Y\in\Sigma^{\infty}$ align over all finite prefixes, their constructive dimensions become equal. From Theorem 6, we get that when this happens, the constructive $\Phi$-dimensions also become equal.
Lemma 13.
For any $x,y\in{\mathbb{X}}$, $A,B\subseteq\mathbb{N}$ and any class of computable Cantor coverings $\Phi$, if for all n, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$, then $\mathrm{cdim}^{A}(x)=\mathrm{cdim}^{B}(y)$ and $\mathrm{cdim}^{A}_{\Phi}(x)=\mathrm{cdim}^{B}_{\Phi}(y)$. Here $X$ and $Y$ are the binary expansions of $x$ and $y$ respectively.
6.4 Faithfulness of Cantor Coverings
Using the Point-to-Set Principle and properties of Kolmogorov complexity, we show that the notions of faithfulness for Cantor coverings at Hausdorff and Constructive levels are equivalent.
We first show that if a class of computable Cantor coverings $\Phi$ is faithful with respect to constructive dimension, then $\Phi$ is also faithful with respect to Hausdorff dimension.
Lemma 14.
For any class of computable Cantor coverings $\Phi$, if for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$, then for all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{\Phi}(\mathcal{F})$.
Proof.
We first show that if $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$ for all $\mathcal{F}\subseteq{\mathbb{X}}$ , then for all $A\subseteq\mathbb{N}$ and $x\in[0,1]\setminus\mathbb{Q}$, $\mathrm{cdim}^{A}(x)=\mathrm{cdim}^{A}_{\Phi}(x)$.
Let $B=\emptyset$. From Theorem 5, we have that for all $x\in[0,1]$, and $A\subseteq\mathbb{N}$, there exists a $y\in[0,1]$, such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K(Y\upharpoonleft n)|=o(n)$. Here $X$ and $Y$ are binary expansions of $x$ and $y$ respectively. Therefore from Lemma 13, we have that $\mathrm{cdim}^{A}(x)=\mathrm{cdim}(y)$ and $\mathrm{cdim}^{A}_{\Phi}(x)=\mathrm{cdim}_{\Phi}(y)$.
Since $\Phi$ is faithful with respect to constructive dimension, we have that $\mathrm{cdim}(y)=\mathrm{cdim}_{\Phi}(y)$. Therefore we have that $\mathrm{cdim}^{A}(x)=\mathrm{cdim}_{\Phi}^{A}(x)$.
Let $\mathcal{F}\subset{\mathbb{X}}$ be arbitrary. From the point-to-set principle for dimension of Cantor Coverings (Corollary 3), we have that
${\mathrm{dim}}_{\Phi}(\mathcal{F})$ $=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}_{\Phi}(x)=\min\limits_{A\subseteq\mathbb{N}}\;\sup\limits_{x\in\mathcal{F}}\;\mathrm{cdim}^{A}(x)={\mathrm{dim}}(\mathcal{F}).$
The last equality follows from the point-to-set principle for Hausdorff dimension (Corollary 1). ∎
To prove the converse, we require the construction of set $\mathcal{I}_{s}$ that contains all points in ${\mathbb{X}}$ having constructive dimension equal to $s$.
Definition 29.
Given $s\in[0,\infty)$, define $\mathcal{I}_{s}=\{x\in{\mathbb{X}}\;|\;\mathrm{cdim}(x)=s\}.$
Lutz and Weihrauch [18] showed that the Hausdorff dimension of $\mathcal{I}_{s}$ is equal to $s$. We provide a simple alternate proof of this using the point-to-set principle.
Lemma 15.
Lutz and Weihrauch [18] ${\mathrm{dim}}(\mathcal{I}_{s})=s.$
Proof.
Since for all $x\in\mathcal{I}_{s}$, by definition $cdim(x)=s$, from Lemma 2, it follows that $\mathrm{cdim}(\mathcal{I}_{s})=s$. From Lemma 1, we have that ${\mathrm{dim}}(\mathcal{I}_{s})\leq\mathrm{cdim}(\mathcal{I}_{s})$, from which it follows that ${\mathrm{dim}}(\mathcal{I}_{s})\leq s$.
Consider any $x\in\mathcal{I}_{s}$. We have that $\mathrm{cdim}(x)=s$. Let $A=\emptyset$. From Theorem 5, we have that for all $B\subseteq\mathbb{N}$, there exists a $y\in[0,1]$ such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$ having $\mathrm{cdim}^{B}(Y)=\mathrm{cdim}(Y)$. Here $X$ and $Y$ are binary expansions of $x$ and $y$ respectively. From Lemma 13, we have that $\mathrm{cdim}^{B}(y)=\mathrm{cdim}^{A}(x)=\mathrm{cdim}(x)=s$.
Therefore, we have that for all $B\subseteq\mathbb{N}$, there exits a $y\in{\mathbb{X}}$ such that $\mathrm{cdim}(y)=\mathrm{cdim}^{B}(y)=s$. Hence $Y\in\mathcal{I}_{s}$ and therefore from the point-to-set principle for Hausdorff dimension (Corollary 1), we have that ${\mathrm{dim}}(\mathcal{I}_{s})=\min\limits_{B\subseteq\mathbb{N}}\sup\limits_{y\in\mathcal{I}_{s}}\mathrm{cdim}^{B}(y)\geq s$. ∎
We now show that if a class of computable Cantor coverings $\Phi$ is faithful with respect to Hausdorff dimension, then $\Phi$ is also faithful with respect to constructive dimension.
Lemma 16.
For any class of computable Cantor coverings $\Phi$, if for all $\mathcal{F}\subseteq{\mathbb{X}}$, ${\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{\Phi}(\mathcal{F})$ , then for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$.
Proof.
Given an $x\in{\mathbb{X}}$ having $\mathrm{cdim}(x)=s$, consider the set $\mathcal{I}_{s}$ from Definition 29. From Lemma 15, we have ${\mathrm{dim}}(\mathcal{I}_{s})=s$. Since $\Phi$ is faithful with respect to Hausdorff dimension, we have that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})=s$.
We now go on to show that for any $x\in{\mathbb{X}}$ having $\mathrm{cdim}=s$ ( or equivalently, for any $x\in\mathcal{I}_{s}$), $\mathrm{cdim}_{\Phi}(x)\leq s$. This along with Lemma 12 shows that $\mathrm{cdim}(x)=\mathrm{cdim}_{\Phi}(x)$. From Lemma 2 and Lemma 10, we have that for all $\mathcal{F}\subseteq{\mathbb{X}}$, $\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F})$.
We first show that for all $B\subseteq\mathbb{N}$, there exists a $y\in\mathcal{I}_{s}$ having $\mathrm{cdim}^{B}_{\Phi}(y)=\mathrm{cdim}_{\Phi}(x)$. Let $A=\emptyset$. From Theorem 5, we have that for all $x\in[0,1]$, and $B\subseteq\mathbb{N}$, there exists a $y\in[0,1]$, such that for all $n\in\mathbb{N}$, $|K^{A}(X\upharpoonleft n)-K^{B}(Y\upharpoonleft n)|=o(n)$. Here $X$ and $Y$ are binary expansions of $x$ and $y$ respectively. Setting $A=\emptyset$ and using Lemma 13, we have that $\mathrm{cdim}_{\Phi}(x)=\mathrm{cdim}^{B}_{\Phi}(y)$.
We now show that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})\geq\mathrm{cdim}_{\Phi}(X)$. From the point-to-set principle for dimension of Cantor Coverings (Corollary 3), we have that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})=\min\limits_{B\subseteq\mathbb{N}}\;\sup\limits_{y\in\mathcal{I}_{s}}\;\mathrm{cdim}^{B}_{\Phi}(y)$. From the argument given above we have that for all $B\subseteq\mathbb{N}$, there exists a $y\in\mathcal{I}_{s}$ having $\mathrm{cdim}^{B}_{\Phi}(y)=\mathrm{cdim}_{\Phi}(x)$. So it follows that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})\geq\mathrm{cdim}_{\Phi}(x)$. Since we have already shown that ${\mathrm{dim}}_{\Phi}(\mathcal{I}_{s})=s$, it follows that $\mathrm{cdim}_{\Phi}(x)\leq s$. ∎
Therefore, we have the following theorem which states that for the classes of Cantor coverings $\Phi$, faithfulness with respect to Hausdorff and Constructive dimensions are equivalent notions.
Theorem 7.
For any class of computable Cantor coverings $\Phi$,
$\forall\mathcal{F}\subseteq{\mathbb{X}}\;;\;{\mathrm{dim}}(\mathcal{F})={\mathrm{dim}}_{\Phi}(\mathcal{F})\iff\forall\mathcal{F}\subseteq{\mathbb{X}}\;;\;\mathrm{cdim}(\mathcal{F})=\mathrm{cdim}_{\Phi}(\mathcal{F}).$
6.5 Log limit condition for faithfulness
Albeverio, Ivanenko, Lebid and Torbin [1] showed that the Cantor coverings $\Phi_{Q}$ generated by the Cantor series expansions of $Q=\{n_{k}\}_{k\in\mathbb{N}}$ can be faithful as well as non faithful with respect to Hausdorff dimension depending on $Q$. Interestingly, they showed that the Hausdorff dimension faithfulness of Cantor coverings can be determined using the terms $n_{k}$ in $Q$.
Theorem 8 (Albeverio, Ivanenko, Lebid and Torbin [1]).
A family of Cantor coverings $\Phi_{Q}$ generated by $Q=\{n_{k}\}_{k\in\mathbb{N}}$ is faithful with respect to Hausdorff dimension if and only if $\lim\limits_{k\rightarrow\infty}\frac{\log n_{k}}{\log n_{1}\cdot n_{2}\dots n_{k-1}}=0.$
Using the above result and the result that faithfulness at the constructive level is equivalent to faithfulness with respect to Hausdorff dimension (Theorem 7), we have that the condition stated above provides the necessary and sufficient conditions for Cantor series coverings to be faithful for constructive dimension.
Theorem 9.
A family of Cantor coverings $\Phi_{Q}$ generated by $Q=\{n_{k}\}_{k\in\mathbb{N}}$ is faithful with respect to constructive dimension if and only if
$\lim_{k\rightarrow\infty}\frac{\log n_{k}}{\log n_{1}.n_{2}\dots n_{k-1}}=0.$ (1)
The Cantor series expansion is a generalization of the base-$b$ representation, which is the special case when $n_{k}=b$ for all $k\in\mathbb{N}$. That is $Q_{b}=\{b\}_{n\in\mathbb{N}}$. Since the condition in Theorem 9 is satisfied by $Q_{b}$ for any $b\in\mathbb{N}$, we have the following result by Hitchcock and Mayordomo about the base invariance of constructive dimension.
Corollary 4 (Hitchcock and Mayordomo [9]).
For any $x\in[0,1]$ and $k,l\in\mathbb{N}\setminus\{1\}$, $\mathrm{cdim}_{(k)}(x)=\mathrm{cdim}_{(l)}(x).$ where $\mathrm{cdim}_{(k)}(x)$ represents the constructive dimension of $x$ with respect to its base-$k$ representation.
Note that condition (1) classifies the Cantor series expansions on the basis of constructive dimension faithfulness. As an example, when $n_{k}={2^{k}}$, condition (1) holds, and therefore $Q=\{2^{k}\}_{k\in\mathbb{N}}$ is faithful for constructive dimension. However, when $n_{k}={2^{2^{k}}}$, condition (1) does not hold, and therefore $Q=\{2^{2^{k}}\}_{k\in\mathbb{N}}$ is not faithful for constructive dimension.
7 Conclusion and Open Problems
We develop a constructive analogue of $\Phi$-dimension and prove a Point-to-Set principle for $\Phi$-dimension. Using this, we show that for Cantor series representations, constructive dimension faithfulness and Hausdorff dimension are equivalent notions. We also provide a loglimit condition for faithfulness of Cantor series expansions.
The following are some problems that remain open
  1. 1.
    Are the faithfulness at constructive and Hausdorff levels equivalent for all computable family of covering sets $\Phi$
  2. 2.
    What is the packing dimension analogue of faithfulness, is there any relationship between faithfulness of Hausdorff dimension and packing dimension ?
  3. 3.
    Is there any relationship between faithfulness of constructive dimension and constructive strong dimension ?